### v1

# CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?

CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?
B J Dalton
Centre for Quantum Science and Technology Theory, Swinburne University of Technology, Melbourne, Victoria 3122, Australia
1 Abstract
In this paper we consider the description by a Bell non-local hidden variable theory of bipartite quantum states with two observables per sub-system. We derive Bell inequalities of the Collins-Gisin.-Liden-Massar-Popescu type which involve combinations of the probabilities of related outcomes for measurements for the four pairs of sub-system observables. It is shown that the corresponding quantum theory expressions violate the Bell inequalities in the case of the maximally entangled state of the bipartitite system. This shows that quantum theory can not be underpinned by a non-local hidden variable theory. So as local hidden variable theory has already been shown to conflict with quantum theory, it follows that quantum theory can not be understood in terms of any hidden variable theory (which Einstein originally had expected).
2 INTRODUCTION
This paper deals with the issue of whether quantum theory needs to be completed or underpinned by a non-local hidden variable theory (NLHVT), an issue that originates from the 1930’s papers of Einstein [1] and Schrodinger ([2], [3]) and the 1960’s paper of Bell ([4]). The key word is non-local,as it is now widely accepted that quantum theory is not underpinned by a local hidden variable theory.
2.1 Quantum Paradoxes
The 1930’s papers identified two paradoxical features arizing from the Copenhagen interpretation of quantum theory (QTHY), even for the simplest case of bipartite systems. The first was macroscopic entanglement – such as in the Schrodinger cat experiment, wherein states could exist in bipartite systems in which a sub-system could simultaneously occupy two macroscopically distinct states. Here the system state was an entangled state of a cat and a radioactive atom, involving both the cat being dead and the cat being alive. This conflicted with common sense - the cat should be either dead or alive. The second was EPR steering - measuring the value for an observable chosen for one sub-system $A$ sometimes could instantaneously affect the outcome for measuring the value of an observable in a second sub-system $B$ that was spatially well-separated. This contradicted relativity and for a system involving two particles in a state with a well-defined position difference, also led to a violation of the Heisenberg U Principle in terms of evidently being able to assign precise values to the position and momentum of one of the two particles. So maybe the particle really does have both a definite position and momentum - and the cat is either dead or alive, whilst the probababilistic feature of quantum theory is just similar to that in classical statistical mechanics ?
2.2 Quantum Theory Incompleteness ?
Thus arose Einstein’s question: [1]. Does quantum theory (though agreeing with experiment) require completion via an underlying theory that determines the real values of physical quantities - such as the position and momentum (for the particle) or being alive or being dead (for the cat), even if these are hidden within the probability distribution ?  Such underlying theories are generally referred to as hidden variable theories.(HVT ).
2.3 Local Hidden Variable Theory and Bell Non-Locality
In response, Bell introduced [4]  a so-called local hidden variable theory (LHVT) for describing the quantum states of composite systems. These states are referred to as Bell local states. In the original deterministic version the hidden variables (HV) determine the actual outcome for each separate sub-system, when any of its observables are measured.  In the general version of LHVT the HV essentially act like classical phase space variables, and determine the probabilities of outcomes for each sub-system separately when any of its observables are measured (the locality feature). A key idea is that the HV only depend on state preparation process -- not on the subsequent choice of what observables will be measured.
LHVT provides predictions for inequalities involving results based on mean values or the probabilities of joint measurement outcomes of pairs of observables from the two sub-systems - generically referred to as Bell inequalities. We can then test whether the inequality predictions based on LHVT are consistent both with quantum theory predictions and with experiment. As described below, some Bell inequality violations show LHVT fails. So does this mean that quantum theory could be underpinned or completed via a non-local HVT (NLHVT) ?
In this paper a non-local hidden variable theory will be presented, and Bell inequalities obtained based on this NLHVT. These inequalities are of the Collins, Gisin, Linden, Massar Popescu (CGLMP) type (see [5], [6]). It will then be shown that in a maximally entangled state in a bipartite system clear violations of these Bell inequalities occur, thus showing that it is not posible to underpin or complete quantum theory via a non-local hidden variable theory. As well as being an important quantum foundations issue, there are also implications also for quantum technology, since systems for secure communications, quantum information processing, etc depend on there being different outcomes predicted for quantum theory and for hidden variable theory.
It should be noted that previous to the work of Bell [4], a different kind of non-local hidden variable theory was introduced by Bohm [7] in the 1950’s, based on earlier work by de Broglie . The approach involves deterministic particle trajectories dependent on a so-called quantum  potential determined (in the non-relativistic version) from the multi-particle Schrodinger wave function for the whole universe. The particle positions act like hidden variables and non-locality refers to the feature that the velocity of any one particle depends on the positions of all the other particles, no matter where their relative locations may be. Interest in Bohmian mechanics and its extensions still continues [8].
2.4 Plan of Paper
First however, we need to outline the key equations describing both LHVT and NLHVT and present a well-known example of a Bell inequality that is inconsistent with quantum theory. Next we point out how Bell inequalities are also relevant to discussing quantum entanglement and EPR steering based on  a classification of Bell local states into three categories involving so-called local hidden quantum states [9]. The material outlining this is in Appendix A. We then introduce the so-called CGLMP Bell inequalities [5] and set out the relevant equations for the inequalities involving the quantity $S$, the observables and the test quantum state, both for the LHVT and NLHVT cases. The next section covers the basic matrix equation method for determining the requirements for possible Bell inequalities of the form $S\leq 4,S\leq 3,S\leq 2,S\leq 1$ in the NLHVT situation  The Bell test requiremennts in each of these cases are set out, with the detailed derivations being presented in Appendix B. The last part of this section then sets out the quantum theory expressions for $S$. A comparison of the quantum theory numerical results for $S$ with the NLHVT results then follows. The paper ends with a Summary and Conclusions.
3 QTHY & HVT BASIC EQUATIONS
QTHY and LHVT joint probabilities $P(\alpha,\beta|\Omega_{A},\Omega_{B})$ for measurements of observables $\Omega_{A}$, $\Omega_{B}$ for sub-systems $A$, $B$ with outcomes $\alpha$, $\beta$ are:
$P(\alpha,\beta|\Omega_{A},\Omega_{B},\rho)_{Q}$ $=$ $Tr((\widehat{\Pi}_{\alpha}^{A}\otimes\widehat{\Pi}_{\beta}^{B})\widehat{\rho})$ (1)
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (2)
For QTHY $\widehat{\Pi}_{\alpha}^{A}$ is the projector onto the space of eigenvectors for operator $\widehat{\Omega}_{A}$ with eigenvalue $\alpha$. $\widehat{\rho}$ is density opr specifying the preparation process $c$ of quantum state. For LHVT, hidden variables $\lambda$ determined with probability $P(\lambda|c)$ for preparation process $c$.
In the probabilistic form of LHVT the hidden variables determine each local sub-system measurement probabilities  $P(\alpha|\Omega_{A},\lambda)$, $P(\beta|\Omega{}_{B},\lambda)$ for classical observables $\Omega_{A}$, $\Omega_{B}$ which  are combined using classical probability theory to determine joint measurement probabilities. Since the probabilities for all outcomes must add up to unity we have $\sum_{\alpha}P(\alpha|\Omega_{A},\lambda)=1$  and $\sum_{\beta}P(\beta|\Omega_{B},\lambda)=1$.
In the probabilistic form of NLHVT the hidden variables determine the combined sub-system measurement probabilities  $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$, for classical observables $\Omega_{A}$, $\Omega_{B}$  In the non–local hidden variable theory there is no separate (or local) HVT probability for each sub-system.
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{NLHVT}=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$ (3)
Again, the sum of the probabilities for all outcomes must be unity,so $\sum_{\alpha,\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)=1$.
3.1 Joint Measurement Mean Values
QTHY and LHVT expressions for the mean values of joint measurement outcomes for $\Omega_{A},\Omega_{B}$ are:
$\left\langle\Omega_{A}\otimes\Omega_{B}\right\rangle_{Q}$ $=$ $Tr(\widehat{\Omega}_{A}\otimes\widehat{\Omega}_{B})\widehat{\rho}$ (4)
$\left\langle\Omega_{A}\otimes\Omega_{B}\right\rangle_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\langle\Omega_{A}(\lambda)\right\rangle\left\langle\Omega_{B}(\lambda)\right\rangle$ (5)
where $\left\langle\Omega_{A}(\lambda)\right\rangle=\sum_{\alpha}\alpha P(\alpha|\Omega_{A},\lambda)$ is LHVT mean value for measurement of $\Omega_{A}$ when hidden variables are $\lambda$.
The NLHVT expression for the mean values of joint measurement outcomes for $\Omega_{A},\Omega_{B}$ is
$\left\langle\Omega_{A}\otimes\Omega_{B}\right\rangle_{NLHVT}=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\langle(\Omega_{A}\otimes\Omega_{B})(\lambda)\right\rangle$ (6)
where $\left\langle(\Omega_{A}\otimes\Omega_{B})(\lambda)\right\rangle=\sum_{\alpha,\beta}\alpha\beta\,P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$ is the NLHVT mean value for measurement of $\Omega_{A}$ and $\Omega_{B}$ when hidden variables are $\lambda$.
3.2 Bell Inequality Test for Bell-Locality
The well-known CHSH Bell inequality ([10]) for bipartite system where outcomes $\alpha,\beta$ are restricted to $(+1,-1)$ and involving two measurement choices for each sub-system is
$S$ $=$ $|\left\langle\Omega_{A1}\otimes\Omega_{B1}\right\rangle_{LHVT}+\left\langle\Omega_{A1}\otimes\Omega_{B2}\right\rangle_{LHVT}+\left\langle\Omega_{A2}\otimes\Omega_{B1}\right\rangle_{LHVT}-\left\langle\Omega_{A2}\otimes\Omega_{B2}\right\rangle_{LHVT}\,|\,\leq 2$
This is violated for singlet state of two spin $1/2$ systems (spin operators in units /2Planck-constant-over-$\hbar/2$).
$\left\lfloor\Psi\right\rangle=\left(\left|+1\right\rangle_{S_{\mu}^{A}}\left|-1\right\rangle_{S_{\mu}^{B}}-\left|-1\right\rangle_{S_{\mu}^{A}}\left|+1\right\rangle_{S_{\mu}^{B}}\right)/\sqrt{2}\qquad\mu=x,y,or\;z$ (8)
for $\Omega_{A1}=S_{z}^{A}$, $\Omega_{A2}=S_{x}^{A}$, $\Omega_{B1}=-(S_{x}^{B}+S_{z}^{B})/\sqrt{2}$, $\Omega_{B2}=+(S_{x}^{B}-S_{z}^{B})/\sqrt{2}$, find $S=2\sqrt{2}$.
3.3 Local Hidden Variable Theory Refuted
Thus in some Bell inequalities (such as the CHSH inequalities [10]) quantum states were found in microscopic bipartite systems where failure of Bell LHVT was predicted, and experiments ([11], [12]) agreed with quantum theory. This was sufficient to rule out LHVT accounting for quantum theory in general.
3.4 Quantum Entanglement and EPR Steering ?
Bell locality violation also implies that EPR steering and quantum entanglement both occur, however quantum entanglement or EPR steering can also occur for some Bell local states ([13], [9], [14], [15], [16]). Although this paper is focused on tests for Bell locality violation, Bell tests for quantum entanglement or EPR steering are also important ([16]). For completeness, a brief outline of how LHVT Bell states can be divided into three categories with different featues for Quantum Entanglement and EPR Steering is set out in Appendix A. The violation of other Bell inequalities for LHVT involving local hidden quantum states (LHS) [9] show whether or not Quantum entanglement or EPR steering is occuring.
Bell inequalities also exist for multi-partite systems -- GHZ states [17] or for measurements at three different times -- Leggett-Garg inequalities ([18]).  A recent review dealing with Bell correlations in macroscopic and mesoscopic systems is in Ref.[19]
In the present paper we will be focused on the so-called CGLMP Bell inequatities ([5], [6]) for bipartite systems.
4 CGLMP BELL INEQUALITIES
The CGLMP Bell inequalities will now be discussed - first for the local hidden variable theory case and then for  the non-local hidden variable theory case.
4.1 LHVT Case
A bipartite system with two observables per sub-system, based on HVT joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})\equiv C(j,k,l,m)$ for all four observables $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ ([5], [6]) is considered. To shorten the notation we may write $\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}\equiv j,k,l,m$. In this bipartite case there are the same number $d$ of outcomes for each observable listed $10,1,...,d-$ 1 for each observable $\ j,k,l,m$..
The joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ applies to the overall result of four different sets of measurements, each involving one observable for each sub-system. These measurement choices are ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B2}$) and ($\Omega_{A1},\Omega_{B2}$). The LHVT expression is given by
$P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ (9)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)$
The probability for joint outcomes of one observable for each sub-system satisfy LHVT conditions [20]. A typical case is $P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})$ $=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j}|\Omega_{A1},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)$.
[…]
$P(\Omega_{A1}{\small=}\Omega_{B1})+P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}1})+P(\Omega_{A2}{\small=}\Omega_{B2})+P(\Omega_{B2}{\small=}\Omega_{A1})\leq 3$ (10)
This Bell inequality based on the LHVT expression is violated for maximally entangled state
$\left|\Psi\right\rangle=\sum_{j=0}^{d-1}\left|j\right\rangle_{A}\otimes\left|j\right\rangle_{B}$ (11)
for observables whose eigenstates for eigenvalues $k$, $l$ in the case of observables $\Omega_{Aa}$ , $\Omega_{Bb}$ ($a,b=1,2$) are:
[…]
Here the $\left|j\right\rangle_{A}$ and $\left|j\right\rangle_{B}$ are orthonormal basis states for sub-systems $A$, $B$ respectively. Note that the observables $\Omega_{A1}$, $\Omega_{A2}$ are not assumed to commute, neither are $\Omega_{B1}$, $\Omega_{B2}$. These observables are alo not required to  have any obvious physical meaning. However the eigenvectors $\left|k\right\rangle_{A,a}$, $\left|l\right\rangle_{B,b}$ can be seen to satisfy the expected orthonormality conditions for different $k$, $l$. Thus
$\left\langle k\;|k^{\ast}\right\rangle_{A,a}$ $=$ $\frac{1}{d}\sum_{j=0}^{d-1}\exp i\frac{2\pi}{d}j\left(-k+k^{\ast}\right)$ (14)
$=$ $\frac{1}{d}\frac{(1-\exp i\frac{2\pi}{d}d(-k+k^{\ast}))}{(1-\exp i\frac{2\pi}{d}(-k+k^{\ast}))}=0\qquad if\text{ }k\neq k^{\ast}$
$=$ $1\qquad if\text{ }k=k^{\ast}$
with a similar result for $\left\langle l\;|l^{\ast}\right\rangle_{B,b\text{ .}}$ell non-locality for above CGLMP Bell inequalitity occurs for $d=2$.
The basic quantum expressions for the sub-system probabilities in the case of the maximally entangled state is (see [5])
$P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ $=$ $|(\left\langle k\;|_{A,a}\otimes\right\langle l\;|_{B,b}|)\;\left|\Psi\right\rangle|^{2}$ (15)
$=$ $\frac{1}{d}\;|\,\sum_{j=0}^{d-1}\left\langle k\;|_{A,a}|\;j\right\rangle_{A}\;|\left\langle l\;|_{B,b}|\;j\right\rangle_{B}|^{2}$
$=$ $\frac{1}{d^{3}}|\;\sum_{j=0}^{d-1}\exp-i\frac{2\pi}{d}j(k+\theta_{a})\;\exp-\frac{2\pi}{d}j(-l+\phi_{b})|^{2}$
$=$ $\frac{1}{d^{3}}\;\frac{|(1-\exp(-i2\pi(k-l+\theta_{A}+\phi_{B}))|^{2}}{|(1-\exp(-i\frac{2\pi}{d}(k-l+\theta_{A}+\phi_{B}))|^{2}}$
$=$ $\frac{1}{d^{3}}\;\frac{\sin^{2}(\pi(k-l+\theta_{a}+\phi_{b}))}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{a}+\phi_{b}))}$
$=$ $\frac{1}{2d^{3}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{a}+\phi_{b}))}$
where we have used $\sin^{2}(\pi(k-l+\theta_{a}+\phi_{b}))=\{\sin\pi(k-l)\cos\pi(\theta_{a}+\phi_{b})+(\cos\pi(k-l)\sin\pi(\theta_{a}+\phi_{b})\}^{2}=\{0+(\pm 1)(\pm 1/\sqrt{2})\}^{2}$, noting that the only $\theta_{a}+\phi_{b}$ values are $1/4,-1/4,3/4,1/4$ and $\sin\pi(1/4)=-\sin\pi(-1/4)=\sin\pi(3/4)=1/\sqrt{2}$.
Macroscopic Bell non-locality occurs in other CGLMP Bell inequalities for $d$ large. Experimental verification up to $d=12$ was found by [21].
4.2 NLHVT Case
We will consider a simple CGLMP Bell inequality involving
$S=P(\Omega_{A1}{\small=}\overline{\Omega_{B1}+\Delta_{11}})+P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}\Delta_{12}})+P(\Omega_{A2}{\small=}\overline{\Omega_{B2}+\Delta_{22}})+P(\Omega_{B2}{\small=}\overline{\Omega_{A1}+\Delta_{21})}$ (16)
where the $\Delta_{ij}$ are all integers and it is understood that $+\Delta_{ij}$ means $+\Delta_{ij}(\mathop{\mathrm{m}od}d)$.
As in the LHVT case, there are two observables $\Omega_{A1},\Omega_{A2}$ for sub-system $A$ and two observables $\Omega_{B1},\Omega_{B2}$ for sub-system $B$. The eigenvectors for all the observables are as for the LHVT case. We will also consider quantum probabilities for the same maximally entangled state that was treated in the LHVT case.
First we construct the NLHVT expression for $S$. We have
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,\delta_{j,l+\Delta_{11}}+\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,\delta_{l,k+\Delta_{12}}\\ +\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\,\delta_{k,m+\Delta_{22}}+\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\,\delta_{l,k+\Delta_{21}}\end{array}\right\}$ (19)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,\delta_{j,l+\Delta_{11}}\,\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\\ +\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,\delta_{l,k+\Delta_{12}}\,\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\\ +\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\,\delta_{k,m+\Delta_{22}}\,\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\\ +\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\,\delta_{m,j+\Delta_{21}}\,\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\end{array}\right\}$ (24)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(\delta_{j,l+\Delta_{11}}+\delta_{k,m+\Delta_{22}}\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(\delta_{l,k+\Delta_{12}}+\delta_{m,j+\Delta 1}\right)\end{array}\right\}$ (27)
The correspondiing expression in LHVT is
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)\\ \times(\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta 1})\end{array}\right\}$ (31)
5 DETERMINING NLHVT BELL INEQUALITIES
First we consider the conditions involving the Kroneka deltas involved in the simple CGLMP inequality and the matrix approach used to treat them. .
5.1 Conditions and Basic Matrix Approach
To derive these inequalities we must consider the conditions where the various Kronecka deltas are equal to unity. These conditions are
$C1$ ::: $\qquad j=l+\Delta_{11}$
$C2$ ::: $\qquad l=k+\Delta_{12}$
$C3$ ::: $\qquad k=m+\Delta_{22}$
$C4$ ::: $\qquad m=j+\Delta_{21}$ (33)
We will consider cases where; Case 1 - all of $C1,C2,C3,C4$ are true (T); Case 2 - three of $C1,C2,C3,C4$ are true and one is false (F); Case 3 - two of $C1,C2,C3,C4$ are true and two are false; Case 4 - one of $C1,C2,C3,C4$ is true and three are false. For these various cases the values of $\delta_{11,22}=\left(\delta_{j,l+\Delta_{11}}+\delta_{k,m+\Delta_{22}}\right)$ and $\delta_{12,21}=\left(\delta_{l,k+\Delta_{12}}+\delta_{m,j+\Delta 1}\right)$ are determined in constructing the Bell inequality.
It turns out that we must consider matrix equations of the form
$M\;\mathbf{v=\Delta}$ (34)
where $M$ is a square matrix and $\mathbf{v}$ is a column vector whose elements are some (or all) of the integers $j,k,l,m$ and $\mathbf{\Delta}$ is a column vector whose elements are sums of the various $\Delta_{ij}$. We will use the result that if the determinent of $M$ is zero ($|M|=0$) then there is only a possible solution for $\mathbf{v}$ if all the elements of the corresponding $\mathbf{\Delta}$ are zero. This will then lead to a set of requirements to lead to Bell inequalities: $S=4$, $S\leq 3$, $S\leq 2$, $S\leq 1$ for the Cases 1, 2, 3, 4.
5.2 Requirements for Bell Inequality Tests
These are derived in Appendix B. The resulting requirements are as follows.
5.2.1 Case1 - S4absent4$\leq$ 4≤ 4
The requirements are
$\Delta_{11}=0\quad\Delta_{12}=0\quad\Delta_{22}=0\quad\Delta_{21}=0$ (35)
If all these requirements are satisfied the Bell inequality $S\leq 4$ would be guaranteed. Note that if these requirements are satisfied then those for $S\leq 3,S\leq 2,S\leq 1$ are not.
5.2.2 Case 2 - S $\leq 3$
The requirements are
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (36)
If this requirement is satisfied the Bell inequality $S\leq 3$ would be guaranteed. Note that if this requirement is satisfied then that for $S\leq 4$ is not.  However those for $S\leq 2,S\leq 1$ could still be satisfied.
5.2.3 Case 3 - S $\leq 2$
The requirements are
$\Delta_{22}$$\neq$ $0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{22}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }b$
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{22}+\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0\qquad\qquad Sub-Case\text{ }e$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0\qquad Sub-Case\text{ }f$
Since we wish to guarantee that the Bell inequality $S\leq 2$ would occur, we can be certain of this by requiring that all the distinct inequalities in the last equation are satisfied. Accordingly, the $And/Or$ are replaced by $And$ and overall we have
$\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{12}\neq 0\quad And\quad\Delta_{21}\neq 0\quad And\quad\Delta_{22}\neq 0$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{21}+\Delta_{11}\neq 0\quad\Delta_{12}+\Delta_{22}\neq 0\quad And\quad\Delta{}_{21}+\Delta_{22}\neq 0$
$\Delta_{11}+\Delta_{21}+\Delta_{22}$$\neq$ $0\quad And\quad\Delta_{22}+\Delta_{12}+\Delta_{11}\neq 0$
$\Delta_{12}+\Delta_{11}+\Delta_{21}$$\neq$ $0\quad And\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0$
Note that if these requirements are satisfied then that for $S\leq 3$ could still be satisfied. For example as $\Delta_{12}+\Delta_{11}\neq 0$ and $\Delta_{21}+\Delta_{22}\neq 0$, we still could have $\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ which is the $S\leq 3$ requirement. However,  those for $S\leq 4$ are not consistent with those here for $S$ $\leq 2$..
5.2.4 Case 4 - S $\leq 1$
The requirements are
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{11}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }b$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{12}\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
Since we wish to guarantee that the Bell inequality $S\leq 1$ would occur, we can be certain of this by requiring that all the distinct inequalities in the last equation are satisfied. Accordingly, the $And/Or$ are replaced by $And$ and overall we have
$\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{12}\neq 0\quad And\quad\Delta_{21}\neq 0\quad And\quad\Delta_{22}\neq 0$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And\quad\Delta_{21}+\Delta_{11}\neq 0\quad\Delta_{12}+\Delta_{22}\neq 0\quad And\quad\Delta_{21}+\Delta_{22}\neq 0$
Note that if these requirements are satisfied then those for $S\leq 2,S\leq 3$ could still be satisfied. For example as $\Delta_{12}+\Delta_{11}\neq 0$ and $\Delta_{21}\neq 0$, we still could have $\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0$ which is a $S\leq 2$ requirement. Similarly if $\Delta_{12}+\Delta_{11}\neq 0$ and $\Delta_{21}+\Delta_{22}\neq 0$ we could still have $\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$, which is the $S\leq 3$ requirement. All the other $S\leq 2$ could be obtained similarly. However those for $S\leq 4$ are not consistent with those here for S $\leq 1$. .
5.3 Quantum Theory Test
Here we compare the quantum theory expressions for $S$ with those from NLHVT based on a specic choice for the $\Delta_{ij}$.
In QTHY the result for the quantity $S$ is given by
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{jl}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\rho)_{Q}\;\delta_{j,l+\Delta_{11}}+\mathop{\displaystyle\sum}\limits_{kl}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\rho)_{Q}\;\delta_{l,k+\Delta_{12}}$
$+\mathop{\displaystyle\sum}\limits_{km}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\rho)_{Q}\;\delta_{k,m+\Delta_{22}}+\mathop{\displaystyle\sum}\limits_{jm}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\rho)_{Q}\;\delta_{m,j+\Delta_{21}}$
and using the QTHY result for $P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ in the case of the maximally entangled state we obtain
$S$ $=$ $\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jl}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-l+\theta_{1}+\phi_{1}))}\;\delta_{j,l+\Delta_{11}}+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{kl}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{2}+\phi_{1}))}\;\delta_{l,k+\Delta_{12}}$ (42)
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{km}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-m+\theta_{2}+\phi_{2}))}\;\delta_{k,m+\Delta_{22}}+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jm}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-m+\theta_{1}+\phi_{2}))}\;\delta_{m,j+\Delta_{21}}$
$=$ $\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{l}\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{11}+\theta_{1}+\phi_{1}))}\;+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{l}\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{12}+\theta_{2}+\phi_{1}))}\;$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{m}\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{22}+\theta_{2}+\phi_{2}))}\;+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{m}\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{21}+\theta_{1}+\phi_{2}))}\;$
$=$ $\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{11}+\theta_{1}+\phi_{1}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{12}+\theta_{2}+\phi_{1}))}$
$+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{22}+\theta_{2}+\phi_{2}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{21}+\theta_{1}+\phi_{2}))}$
Note that adding an integer multiple of $d$ to the $\Delta_{ij}$ does not alter the result.
We could choose $\Delta_{ij}$ in the form $\Delta_{ij}=d\times\Delta{}_{ij}^{\ast}$ where the $\Delta_{ij}^{\ast}$ are all integers. In this case, using $\sin^{2}(\pi\Delta_{ij}^{\ast}+\eta)=(\sin(\pi\Delta_{ij}^{\ast})\cos(\eta)+\cos(\pi\Delta_{ij}^{\ast})\sin(\eta))^{2}=\sin^{2}(\eta)$ we finally obtain
$S$ $=$ $\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(d\Delta_{11}^{\ast}+\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-d\Delta_{12}^{\ast}+\frac{3}{4}))}$ (43)
$+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(d\Delta_{22}^{\ast}+\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-d\Delta_{21}^{\ast}-\frac{1}{4}))}$
$=$ $\frac{1}{2d^{2}}\;\left(\frac{1}{\sin^{2}(\frac{\pi}{4d})}+\frac{1}{\sin^{2}(\frac{3\pi}{4d})}+\frac{1}{\sin^{2}(\frac{\pi}{4d})}+\frac{1}{\sin^{2}(\frac{\pi}{4d})}\right)$
$=$ $\frac{1}{2d^{2}}\;\left(\frac{3}{\sin^{2}(\frac{\pi}{4d})}+\frac{1}{\sin^{2}(\frac{3\pi}{4d})}\right)$
This last result shows that in quantum theory the value of $S$ is independent of the $\Delta_{ij}^{\ast}$. It only depends on $d$. The NLHVT Bell inequalities for $S$ do of course depend on the $\Delta_{ij}^{\ast}$.
For the choice $d=10$ we find that $S\approx 2.52$, so this exceeds $S\leq 2$ and $S\leq 1$ though not $S\leq 4$ and $S\leq 3$ that could occur in NLHVT depending on the requirements for these cases being met. For the choice $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-2,1,-2)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3$ are met, as are the reqirements for $S\leq 1$, though not the requirements for $S\leq 2$. Other choices are also be made. For example with $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-3,1,-3)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3,S\leq 2$ and $S\leq 1$ are all met. Thus we have found cases for $d=10$ where the quantum theory prediction that $S=2.52$ is inconsistent with the NLHVT Bell inequality prediction of $S\leq 2.$
We could also choose $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ as in the original CGLMP Bell inequality. In this case the Bell inequality predicts $S\leq 3$. For this choice the expression for $S$ is given by
$S$ $=$ $\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-1+\frac{3}{4}))}$ (44)
$+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-\frac{1}{4}))}$
$=$ $\frac{2}{d^{2}}\frac{1}{\sin^{2}(\frac{\pi}{4d})}$
For the case of $d=2$ this gives $S=3.41$. For the choice $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ the requirements for $S\leq 4$ are not met, but those for $S\leq 3$ are met,.The requirements for $S\leq 4$, $S\leq 2$ and $S\leq 1$ are not met. However NLHVT is inconsistent with QTHY in this choice of $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ as it was for LHVT where in both  types of HVT the inequality $S\leq 3$ is violated in Q THY.
6 SUMMARY AND CONCLUSION
A brief review of the origins of local hidden variable theory, Bell local states and Bell inequalities has been presented. The classification of Bell local states into three categorise based on the local hidden quantum state approach of Wiseman has been explained. We outline how the Bell inequalities provide tests for whether local hidden variable theory is in conflict with quantum theory and also which categories of Bell local states exhbit quantum entanglement or EPR steering.
The new result in this paper is that a non-local hidden variable theory has been presented, and Bell inequalities obtained based on this NLHVT. These inequalities are of the Collins, Gisin, Linden, Massar Popescu (CGLMP) type.. It has been shown that in a maximally entangled state in a bipartite system clear violations of these Bell inequalities occur, thus showing that it is not posible to underpin or complete quantum theory via a non-local hidden variable theory.
It should be pointed out that the observables chosen have no obvious physical meaning. Also the experimental measurements for the NLHVT Bell inequality terms would difficult to perform. Nor would the maximally entangled state be easy to prepare. However these issues are irrelevant as all that is needed is that the observables are possible quantum observables, the state is a possible quantum state and the Bell ineqquality measurements could be carried out.
The result that a non-local hidden variable theory is also in conflict with quantum theory is of some significance since it shows that quantum theory can not be understood  in terms of hidden variable theory- as Einstein had expected. This is irrespective of whether the hidde variable theory is local or non-local.
7 ACKNOWLEDGEMENTS
The author wishes to acknowledge discussions on quantun foundations with several colleagues - M Reid, S Barnett, B Garraway, L Heaney, J Goold, T Busch, L Rosales, R Teh, B Opanchuk and P Drummond.
8 APPENDIX A - LOCAL HIDDEN QUANTUM STATES, ENTANGLEMENT & EPR STEERING
Quantum states can be divided into entangled and separable states, but in LHVT can also be divided into Bell local and Bell non-local states. These categories overlap [13].
8.1 Separable States
For separable states, prepare sub-systems $A$, $B$ in quantum states $\widehat{\rho}_{R}^{A}$, $\widehat{\rho}_{R}^{B}$, where combined state $\widehat{\rho}_{R}^{A}\otimes\widehat{\rho}_{R}^{B}$ occurs with classical probability $P_{R}$ [13]. With density opr $\widehat{\rho}_{sep}=\sum_{R}P_{R}\;\widehat{\rho}_{R}^{A}\otimes\widehat{\rho}_{R}^{B}$ and $P(\alpha|\Omega_{A},R)=Tr_{A}(\widehat{\Pi}_{\alpha}^{A}\widehat{\rho}_{R}^{A})$ etc.,joint outcome prob is given by a LHVT form:
$P(\alpha,\beta|\Omega_{A},\Omega_{B},\rho_{sep})_{Q}=\sum_{R}P_{R}\,P(\alpha|\Omega_{A},R)\times P(\beta|\Omega_{B},R)$ (45)
This is of the same form as for Bell local states, showing that all separable states are Bell-local ($R\rightarrow\lambda$). Hence Bell non-local states must be entangled.
However, some Bell local states are also entangled  [13]. For Bell-local states which are separable, both sub-systems are associated with a quantum density operator. Hence idea of a local hidden quantum state (LHQS) was introduced [9], [14], [15].
8.2 Local Hidden Quantum State
Bell-local states in bipartite systems can be further divided into three disjoint categories, depending on two, one or none of the sub-systems being associated with a local hidden quantum state [9], [14], [15]. LHQS $\widehat{\rho}_{C}(\lambda)$ are to be possible quantum states -- these must comply with conditions such as the super-selection rules of there being no coherences between states with differing (massive) particle numbers.
8.3 Bell-Local States - Three Categories
For the bipartite case - there are three categories of Bell-local states
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\textstyle\sum}\limits_{\lambda}P_{Q}(\alpha|\Omega_{A},\lambda)\,P_{Q}(\beta|\Omega_{B},\lambda)\,P(\lambda|c)\qquad Cat\;1$
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\textstyle\sum}\limits_{\lambda}P(\alpha|\Omega_{A},\lambda)\,P_{Q}(\beta|\Omega_{B},\lambda)\,P(\lambda|c)\qquad Cat\;2$
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\textstyle\sum}\limits_{\lambda}P(\alpha|\Omega_{A},\lambda)\,P(\beta|\Omega_{B},\lambda)\,P(\lambda|c)\qquad Cat\;3$
where subscript $Q$ indicates sub-system measurement probability determined from a local hidden quantum state $\widehat{\rho}_{C}(\lambda)$, via $P_{Q}(\gamma|\Omega_{C},\lambda)=Tr_{C}(\widehat{\Pi}_{\gamma}^{C}\widehat{\rho}_{C}(\lambda))$. Bell inequalities will differ depending on whether there are sub-systems associated with a LHQS.
8.4 Tests for Quantum Entanglement and EPR Steering
For the case of two modes. we introduce spin operators $S_{x}=\frac{1}{2}(b^{{\dagger}}a+a^{{\dagger}}b),S_{y}=\frac{1}{2i}(b^{{\dagger}}a-a^{{\dagger}}b),S_{z}=\frac{1}{2}(b^{{\dagger}}b-a^{{\dagger}}a)$.
Category 1 states are same as separable states since both sub-system probabilities are given by quantum expressions. Violation of inequalities based on Category 1 states $\rightarrow$ Quantum entanglement (non-separability) occurs. Example: Spin squeezing test for two mode entanglement [22] $\left\langle\Delta S_{x}^{2}\right\rangle<\frac{1}{2}|\left\langle S_{z}\right\rangle$ or $|\left\langle\Delta S_{y}^{2}\right\rangle<\frac{1}{2}|\left\langle S_{z}\right\rangle|$.
For Category 2 states only one sub-system probability ($B$) is given by quantum expression involving LHS quantum density operator -- $\widehat{\rho}^{B}(\lambda)$. Here no EPR steering of $B$ by $A$ occurs $\rightarrow$ Outcomes $\beta$ for $\Omega_{B}$ depend only on $\widehat{\rho}_{B}(\lambda)$- not on $\alpha,\Omega_{A}$. Violation of inequalities based on Category 2 states $\rightarrow$ EPR steering of $B$ outcomes by $A$ measurement occurs. Example: generalised Hillary-Zubairy planar spin variance test for two mode EPR steering [23], [16]. $\left\langle\Delta S_{x}^{2}\right\rangle+\left\langle\Delta S_{y}^{2}\right\rangle<(\frac{1}{4}\left\langle N\right\rangle-\frac{1}{2}\left\langle S_{z}\right\rangle)$.
8.5 Tests for Bell Non-Locality
For Category 3 states neither of sub-system probabilities given by quantum expression - no local hidden quantum state. . Inequality violations based on Category 3 states $\rightarrow$ Bell locality cannot apply. Example: CHSH Bell Inequality (above) for two spin 1/2 systems. Thus Bell non-locality is shown by violations of inequalities based on
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (47)
8.6 Bell - Non-Local States
A fourth category of states are the Bell non-local states.
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{NLHVT}=\mathop{\textstyle\sum}\limits_{\lambda}P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)\,\,P(\lambda|c)\qquad Cat\;4$ (48)
Here there is no separate $P(\gamma|\Omega_{C},\lambda)\,$for each sub-system. Violation of inequalities based on Category 4 states show that non-local hidden variable theories also fail.
8.7 Overall Scheme
The overall scheme can be pictured as (see [24], [16])  Here we have assumed that Bell non-local states are quantum entangled states.
Figure 1: The classifcation of hidden variable Bell local states into three categories, plus the hidden variable Bell non-local states. Also the classification of quantum states as separable and entangled. The features of the various types of states are stated.
9 APPENDIX B - REQUIREMENTS FOR BELL TESTS
In this Appendix  we will derive the requirements for Bell Inequality tests in the NLHVT situation.
9.1 Case 1 - Bell Inequality - All Conditions T
If $C1=T,C2=T,C3=T,C4=T$ we then have
$\left[\begin{array}[]{cccc}1&0&-1&0\\ 0&-1&1&0\\ 0&1&0&-1\\ -1&0&0&1\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\\ l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{12}\\ \Delta_{22}\\ \Delta_{21}\end{array}\right]$ (49)
As $|M|=0$ there is only a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}=0\quad\Delta_{12}=0\quad\Delta_{22}=0\quad\Delta_{21}=0$ (50)
This solution is $j=k=l=m$ and here we have $\delta_{11,22}=2$ and $\delta_{12,21}=2$’  So in general wecan say that if the requirement is satisfied then $\delta_{11,22}\leq 2$ and $\delta_{12,21}\leq 2$.
Hence
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(2\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(2\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(2+2)$
$\leq$ 44 44 (54)
The Bell inequality is then $S\leq 4$.
9.2 Case 2 - Bell Inequality - Three Conditions T, One F
Here there are four sub-cases
9.2.1 Sub-Case a - $C1=T,C2=T,C3=T,C4=F$
Substituting $C1,C2,C3$ into $C4$ we have
$m$ $=$ $j+\Delta_{21}$ (55)
$=$ $l+\Delta_{11}+\Delta_{21}$
$=$ $k+\Delta_{12}+\Delta_{11}+\Delta_{21}$
$=$ $m+\Delta_{22}+\Delta_{12}+\Delta_{11}+\Delta_{21}$
As $C4=F$ it follows that the last equation is false. Hence the following requirement must be satisfied
$\Delta_{22}+\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0$ (56)
[…]
9.2.2 Sub-Case b - $C1=T,C2=T,C3=F,C4=T$
Substituting $C1,C2,C4$ into $C3$ we have
$k$ $=$ $m+\Delta_{22}$ (57)
$=$ $j+\Delta_{21}+\Delta_{22}$
$=$ $l+\Delta_{11}+\Delta_{21}+\Delta_{22}$
$=$ $k+\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}$
As $C3=F$ it follows that the last equation is false. Hence the following requirement must be satisfied
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (58)
This is the same as for Sub_Case a. In this sub-case we have $\delta_{11,22}=1$ and $\delta_{12,21}=2$. So in general wecan say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 2$.
9.2.3 Sub-Case c - $C1=T,C2=F,C3=T,C4=T$
The requirement and the Bell inequality is the same as for Sub-Cases a and b.
9.2.4 Sub-Case d - $C1=F,C2=T,C3=T,C4=T$
The requirement and the Bell inequality is the same as for Sub-Cases a, b and c.
9.2.5 All Sub-Cases of Case 2
Overall, we see that if the requirement is satisfied, we can either say $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 2$ or $\delta_{11,22}\leq 2$ and $\delta_{12,21}\leq 1$. Any of these possibilities occurs when three of $C1,C2,C3,C4$ are true and one is false.
Hence applying the first possibility
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(1\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(2\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(1+2)$
$\leq$ 33 33 (62)
The second possibility gives the same result
From above the following requirement must be satisfied for Case 2 to cover all sub-cases.
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (63)
The Bell inequality is then $S\leq 3$.
9.3 Case 3 - Bell Inequality - Two Conditions T, Two Conditions F
Here there are six sub-cases
9.3.1 Sub-Case a $C1=T,C2=T,C3=F,C4=F$
Substituting $C1,C2$ into $C3,$ $C4$ we have
$k$ $=$ $m+\Delta_{22}$
$m$ $=$ $j+\Delta_{21}$ (64)
$=$ $l+\Delta_{11}+\Delta_{21}$
$=$ $k+\Delta_{12}+\Delta_{11}+\Delta_{21}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}k\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{22}\\ \Delta_{12}+\Delta_{11}+\Delta_{21}\end{array}\right]$ (65)
Since $C3=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{22}\neq 0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0$ (66)
Since in this sub-case we have $C1=T,C2=T,C3=F,C4=F$we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
9.3.2 Sub-Case b $C1=T,C2=F,C3=T,C4=F$
Substituting $C1,C3$ into $C2,$ $C4$ we have
$l$ $=$ $k+\Delta_{12}$
$=$ $m+\Delta_{22}+\Delta_{12}$
$m$ $=$ $j+\Delta_{21}$ (67)
$=$ $l+\Delta_{11}+\Delta_{21}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{22}+\Delta_{12}\\ \Delta_{11}+\Delta_{21}\end{array}\right]$ (68)
Since $C2=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0$ (69)
Since in this sub-case we have $C1=T,C2=F,C3=T,C4=F$ we have $\delta_{11,22}=2$ and $\delta_{12,21}=0$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 2$ and $\delta_{12,21}=0$.
9.3.3 Sub-Case c $C1=T,C2=F,C3=F,C4=T$
Substituting $C1,C4$ into $C2,$ $C3$ we have
$l$ $=$ $k+\Delta_{12}$
$k$ $=$ $m+\Delta_{22}$ (70)
$=$ $j+\Delta_{21}+\Delta_{22}$
$=$ $l+\Delta_{11}+\Delta_{21}+\Delta_{22}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}-1&+1\\ +1&-1\end{array}\right]\;\left[\begin{array}[]{c}k\\ l\end{array}\right]=\left[\begin{array}[]{c}\Delta_{12}\\ \Delta_{11}+\Delta_{21}+\Delta_{22}\end{array}\right]$ (71)
Since $C2=F,C3=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{12}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (72)
Since in this sub-case we have $C1=T,C2=F,C3=F,C4=T$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
9.3.4 Sub-Case d $C1=F,C2=T,C3=T,C4=F$
Substituting $C2,C3$ into $C1,$ $C4$ we have
$j$ $=$ $l+\Delta_{11}$
$=$ $k+\Delta_{12}+\Delta_{11}$
$=$ $m+\Delta_{22}+\Delta_{12}+\Delta_{11}$
$m$ $=$ $j+\Delta_{21}$ (73)
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{22}+\Delta_{12}+\Delta_{11}\\ \Delta_{21}\end{array}\right]$ (74)
Since $C1=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{22}+\Delta_{12}+\Delta_{11}\neq 0\quad And/Or\quad\Delta{}_{21}\neq 0$ (75)
Since in this sub-case we have $C1=F,C2=T,C3=T,C4=F$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
9.3.5 Sub-Case e $C1=F,C2=T,C3=F,C4=T$
Substituting $C2,C4$ into $C1,$ $C3$ we have
$j$ $=$ $l+\Delta_{11}$
$=$ $k+\Delta_{12}+\Delta_{11}$
$k$ $=$ $m+\Delta_{22}$ (76)
$=$ $j+\Delta_{21}+\Delta_{22}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\end{array}\right]=\left[\begin{array}[]{c}\Delta_{12}+\Delta_{11}\\ \Delta_{21}+\Delta_{22}\end{array}\right]$ (77)
Since $C1=F,C3=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{12}+\Delta_{11}\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0$ (78)
Since in this sub-case we have $C1=F,C2=T,C3=F,C4=T$ we have $\delta_{11,22}=0$ and $\delta_{12,21}=2$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}=0$ and $\delta_{12,21}\leq 2$.
9.3.6 Sub-Case f $C1=F,C2=F,C3=T,C4=T$
Substituting $C3,C4$ into $C1,$ $C2$ we have
$j$ $=$ $l+\Delta_{11}$
$l$ $=$ $k+\Delta_{12}$ (79)
$=$ $m+\Delta_{22}+\Delta_{12}$
$=$ $j+\Delta_{21}+\Delta_{22}+\Delta_{12}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}-1&+1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}l\\ j\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{21}+\Delta_{22}+\Delta_{12}\end{array}\right]$ (80)
Since $C1=F,C2=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0$ (81)
Since in this sub-case we have $C1=F,C2=F,C3=T,C4=T$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
9.3.7 All Sub-Cases of Case 3
Overall, we see that if the requirement is satisfied, we can either say $\delta_{11,22}=0$ and $\delta_{12,21}\leq 2$ or $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$ or $\delta_{11,22}\leq 2$ and $\delta_{12,21}=0$ Any of these possibilities occurs when two of $C1,C2,C3,C4$ are true and two are false.
Hence applying the second possibility
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(1\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(1\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(1+1)$
$\leq$ 22 22 (85)
The two other possibilities give the same result.
From above the following requirements must be satisfied for Case 3 to cover all sub-cases.
$\Delta_{22}$$\neq$ $0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{22}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }b$
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{22}+\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0\qquad\qquad Sub-Case\text{ }e$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0\qquad Sub-Case\text{ }f$
The Bell inequality is then $S\leq 2$.
9.4 Case 4 Bell Inequality - One Condition T, Three Conditions F
Here there are four sub-cases.
9.4.1 Sub-Case a $C1=T,C2=F,C3=F,C4=F$
Substituting $C1$ into $C2,$ $C3,$ $C4$ we have
$l$ $=$ $k+\Delta_{!2}$
$k$ $=$ $m+\Delta_{22}$
$m$ $=$ $j+\Delta_{21}$ (87)
$=$ $l+\Delta_{11}+\Delta_{21}$
Thus we have the marix equation
$\left[\begin{array}[]{ccc}-1&+1&0\\ +1&0&-1\\ 0&-1&+1\end{array}\right]\;\left[\begin{array}[]{c}k\\ l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{!2}\\ \Delta_{22}\\ \Delta_{11}+\Delta_{21}\end{array}\right]$ (88)
Since $C2=F,C3=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{!2}\neq 0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0$ (89)
Since in this sub-case we have $C1=T,C2=F,C3=F,C4=F$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=0$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}=0$.
9.4.2 Sub-Case b $C1=T,C2=F,C3=F,C4=F$
Substituting $C2$ into $C1,$ $C3,$ $C4$ we have
$j$ $=$ $l+\Delta_{11}$
$=$ $k+\Delta_{!2}+\Delta_{11}$
$k$ $=$ $m+\Delta_{22}$
$m$ $=$ $j+\Delta_{21}$ (90)
Thus we have the marix equation
$\left[\begin{array}[]{ccc}+1&-1&0\\ 0&+1&-1\\ -1&0&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{!2}+\Delta_{11}\\ \Delta_{22}\\ \Delta_{21}\end{array}\right]$ (91)
Since $C1=F,C3=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{!2}+\Delta_{11}\neq 0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{21}\neq 0$ (92)
Since in this sub-case we have $C1=F,C2=T,C3=F,C4=F$ we have $\delta_{11,22}=0$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}=0$ and $\delta_{12,21}\leq 1$.
9.4.3 Sub-Case c $C1=F,C2=F,C3=T,C4=F$
Substituting $C3$ into $C1,$ $C2,$ $C4$ we have
$j$ $=$ $l+\Delta_{11}$
$l$ $=$ $k+\Delta_{12}$
$=$ $m+\Delta_{22}+\Delta_{12}$
$m$ $=$ $j+\Delta_{21}$ (93)
Thus we have the marix equation
$\left[\begin{array}[]{ccc}+1&-1&0\\ 0&+1&-1\\ -1&0&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{22}+\Delta_{12}\\ \Delta_{21}\end{array}\right]$ (94)
Since $C1=F,C2=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}\neq 0$ (95)
Since in this sub-case we have $C1=F,C2=F,C3=T,C4=F$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=0$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}=0$.
9.4.4 Sub-Case d $C1=F,C2=F,C3=F,C4=T$
Substituting $C4$ into $C1,$ $C2,$ $C3$ we have
$j$ $=$ $l+\Delta_{11}$
$l$ $=$ $k+\Delta_{12}$
$k$ $=$ $m+\Delta_{22}$ (96)
$=$ $j+\Delta_{21}+\Delta_{22}$
Thus we have the marix equation
$\left[\begin{array}[]{ccc}+1&0&-1\\ 0&-1&+1\\ -1&+1&0\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\\ l\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{12}\\ \Delta_{21}+\Delta_{22}\end{array}\right]$ (97)
Since $C1=F,C2=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}\neq 0\quad And/Or\quad\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0$ (98)
Since in this sub-case we have $C1=F,C2=F,C3=F,C4=T$ we have $\delta_{11,22}=0$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}=0$ and $\delta_{12,21}\leq 1$.
9.4.5 All Sub-Cases of Case 4
Overall, we see that if the requirement is satisfied, we can either say $\delta_{11,22}=0$ and $\delta_{12,21}\leq 1$ or $\delta_{11,22}\leq 1$ and $\delta_{12,21}=0$. Any of these possibilities occurs when one of $C1,C2,C3,C4$ is true and three are false.
Hence applying the second possibility
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(1\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(0\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(1+0)$
$\leq$ 11 11 (102)
The other possibility gives the same result.
From above the following requirements must be satisfied for Case 4 to cover all sub-cases.
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{11}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }b$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{12}\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
The Bell inequality is then $S\leq 1$.