### v2

# CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?

CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?
B J Dalton
Centre for Quantum Science and Technology Theory, Swinburne University of Technology, Melbourne, Victoria 3122, Australia
0.1 Abstract
In this paper we consider the description by a general Bell-type non-local hidden variable theory of bipartite quantum states with two observables per sub-system. We derive Bell inequalities of the Collins-Gisin.-Liden-Massar-Popescu type which involve combinations of the probabilities of related outcomes for measurements for the four pairs of sub-system observables. It is shown that the corresponding quantum theory expressions violate the Bell inequalities in the case of the maximally entangled state of the bipartitite system. The CHSH Bell inequality is also derived from this general CGLMP Bell-type non-local hidden variable theory. This shows that quantum theory can not be underpinned by a Bell-type non-local hidden variable theory. So as a general Bell-type local hidden variable theory has already been shown to conflict with quantum theory, it follows that quantum theory can not be understood in terms of any CGLMP Bell-type hidden variable theory - local or non-local.
1 INTRODUCTION
This paper deals with the issue of whether quantum theory needs to be completed or underpinned by a non-local hidden variable theory (NLHVT), an issue that originates from the 1930’s papers of Einstein [1] and Schrodinger ([2], [3]) and the 1960’s paper of Bell ([5]). The key word is non-local,as it is now widely accepted that quantum theory is not underpinned by a local hidden variable theory (LHVT).
1.1 Quantum Paradoxes
The 1930’s papers identified two paradoxical features arizing from the Copenhagen interpretation of quantum theory (QTHY), even for the simplest case of bipartite systems. The first was macroscopic entanglement – such as in the Schrodinger cat experiment, wherein states could exist in bipartite systems in which a sub-system could simultaneously occupy two macroscopically distinct states. Here the system state was an entangled state of a cat and a radioactive atom, a quantum superposition state involving both the cat being dead and the cat being alive. This conflicted with common sense - the cat should be either dead or alive. The second was EPR steering - measuring the value for an observable chosen for one sub-system $A$ sometimes could instantaneously affect the outcome for measuring the value of an observable in a second sub-system $B$ that was spatially well-separated. This contradicted relativity and for a system involving two particles in a possible state with a well-defined position difference (and a well-defined momentum sum), also led to a violation of the Heisenberg Uncertainty Principle in terms of evidently being able to assign precise values to both the position and momentum of one of the two particles. So maybe the particle really does have both a definite position and momentum - and the cat is either dead or alive, whilst the probabilistic feature of quantum theory is just similar to that in classical statistical mechanics ?
1.2 Quantum Theory Incompleteness ?
Thus arose Einstein’s question: [1]. Does quantum theory (though agreeing with experiment) require completion via an underlying theory that determines the actual values of physical quantities (realism) - such as the position and momentum (for the particle) or being alive or being dead (for the cat), even if these are hidden within the probability distribution ?  Since the work of Bohm [4] and Bell [5] such underlying (or underpinning) theories are now generally referred to as hidden variable theories.(HVT ).
1.3 Local and Non-Local Hidden Variable Theory
In response, Bell introduced [5]  a so-called local hidden variable theory (LHVT) for describing the quantum states of composite systems. These states are referred to as Bell local states. In the original deterministic version the hidden variables (HV) determine the actual outcome for each separate sub-system, when any of its observables are measured.  In the general version of LHVT the HV essentially act like classical phase space variables, and determine the probabilities of outcomes for each sub-system separately when any of its observables are measured (the locality feature). A key idea is that the HV only depend on state preparation process -- not on the subsequent choice of what observables will be measured.
LHVT provides predictions for inequalities involving results based on mean values or the probabilities of joint measurement outcomes of pairs of observables from the two sub-systems - generically referred to as Bell inequalities. We can then test whether the inequality predictions based on LHVT are consistent both with quantum theory predictions and with experiment. As described below, some Bell inequality violations show LHVT fails. So does this mean that quantum theory should be underpinned or completed via a Bell-type non-local HVT (NLHVT) ?
In this paper a general non-local hidden variable theory will be presented, and Bell inequalities obtained based on this NLHVT.We will show that the famous CHSH [6] Bell inequality can be established from NLHVT as well as from LHVT. We then consider Bell inequalities of the Collins, Gisin, Linden, Massar Popescu (CGLMP) type (see [7], [8]). It will then be shown that in a maximally entangled state in a bipartite system there are clear violations of these Bell inequalities, thus showing that it is not posible to underpin or complete quantum theory via this general Bell-type non-local hidden variable theory. As well as being an important quantum foundations issue, there are also implications also for quantum technology, since systems for secure communications, quantum information processing, etc depend on there being different outcomes predicted for quantum theory and for hidden variable theory.
1.4 Bohmian Mechanics
It should be noted that previous to the work of Bell [5], a different kind of non-local hidden variable theory was introduced by Bohm [4] in the 1950’s, based on earlier work by de Broglie (see [9] for a recent account). The approach involves deterministic particle trajectories dependent on a so-called quantum potential determined (in the non-relativistic version) from the multi-particle Schrodinger wave function for the system. The particle positions act like hidden variables and non-locality refers to the feature that the velocity of any one particle depends on the positions of all the other particles, no matter where their relative locations may be. This would violate special relativity, but as the related version of quantum theory is also non-relativistic this outcome is merely being consistent. As the derivation of the equations for Bohmian mechanics (BM) invokes the existence of the wave-function, assumes the correctness of the time-dependent Schrodinger equation (TDSE) and that the modulus squared of the wave function is the position probability density for the particles, it may be concluded that Bohmian mechanics is based on just the same equations as quantum theory but with some extra assumptions and ontological features (the deterministic trajectories) added. The particle trajectory feature is based on deriving a continuity equation for the position probabilty density and identifying the gradient of the phase $S$ of the wave-function (divided by the partticle mass) with the velocity for the deterministic particle trajectory - whose existence is now incorporated into the theory. This identification is based on a continuity equation for the position probability derived from equations for the phase and amplitude of the wave-function obtained from the TDSE. So the addition of deterministic trajectories complies with Einstein’s suggestion of completing quantum theory by adding new features - the deterministic trajectories. Whether the particle positions should be regarded as ”hidden variables” is a moot point. Unlike the hidden variables $\lambda$ introduced by Bell [5] (see below), the particle positions could in principle be observed, and they obey well defined Hamilton-Jacobi equations involving the quantum potential - in contrast to the unspecified equations for the $P(\lambda|c)$ and $P(\alpha|\Omega_{A},\lambda)$ in Bell’s approach. BM is rather like classical  physics and uses a more familiar mathematics based on calculus rather than the mathematics of linear vector spaces and Hermitian  operators. Many of the characteristic results in standard quantum theory can also be shown in BM. For example the component $L_{z}$ of the orbital  angular momentum of a particle is $xp_{y}-yp_{x}$, which becomes $x(\partial S/\partial y)-y(\partial S/\partial x)$ $=(\partial S/\partial\phi)$ from the particle trajectory assumption in which $p_{y}=mv_{y}$ etc.. As the stationary solution of the Schrodinger equation for a spherical potential involves a phase $S=\beta-Et+\hslash m\phi$ ($m$ an integer) it follows that $L_{z}$   has a quantised value $m\hslash$. Other quantum theory features such as the symmetry or anti-symmetry of the wave-function for identical particles are just assumed true in Bohmian mechanics. Relativistic versions of Bohmian mechanics are still being developed, as are quantum field theory versions. As far as is known, Bohmian mechanics does not predict any new physical results, but it has been applied to the well-known two-slit experiment for single particles [10] - where it is shown that the particle trajectories are affected by the quantum potential (that reflects the interference effects on the wave-function due to the two slits), and this then leads to the periodic particle position probability distribution on the observation screen. Whether Bohmian mechanics can describe bipartite systems for all pairs of non-commuting observables in each sub-system is an issue that is not examined in the present paper. However, interest in Bohmian mechanics and its extensions still continues [11].
1.5 Other Proposed Non Local HVT
Another approach to completing quantum theory is the Many Worlds Interpretation (MWI) [12]. Essentially the probabilistic feature of quantum theory - where in general each possible outcome of measuring an observable only occurs with a certain probability - is replaced with a deterministic picture in which all the different outcomes actually occur, but now in dfferent universes. The MWI has recently been discussed by Vaidman [13]. Leggett [14], introduced a so-called Crypto-Nonlocal (CN) hidden variable theory.In the context of the present paper, the LHVT expression (2) $P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ for the bipartite system case would be replaced in the CN theory by $P(\alpha|\Omega_{A},\Omega_{B},\lambda)P(\beta|\Omega_{A},\Omega_{B},\lambda)$, so that although the outcome for measuring either sub-system observable does not depend on the outcome for measuring the other sub-system observable, it does depend on the choice of the other observable. It is not clear whether this would produce different Bell inequalities from the NLHVT replacement $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$. CN theory has been found to violate quantum theory and experiment [15].
1.6 Plan of Paper
First however, we need to outline the key equations describing both LHVT and a form of NLHVT in the simplest case of bipartite systems. Here the basic probability involves the outcomes for one observable in each sub-system. We then discuss the well-known CHSH example of a Bell inequality based on LHVT that is inconsistent with quantum theory. We also outline the idea behind LHVT. For reasons of completeness we next point out how Bell inequalities are also relevant to discussing quantum entanglement and EPR steering based on  a classification of Bell local states into three categories involving so-called local hidden quantum states [16]. The material outlining this is in Appendix A. We then introduce the so-called CGLMP Bell inequalities [7]. Here the basic probability involves the outcomes for two observables in each sub-system, resulting in more general LHVT andNLHVT forms. We then show a novel result - namely that the CHSH Bell inequality can also be proved for this more general NLHVT. The violation of the CHSH Bell inequality for this NLHVT approach proves that quantum theory cannot be underpinned by this type of NLHVT. The CGLMP inequalities themselves are applied in a bipartite system where each sub-system has two observables, each with the same number $d$ of outcomes) and set out the relevant equations for the inequalities - involving the quantity $S$, and introduce the observables and the test quantum state, both for the LHVT and NLHVT cases. The quantity $S$ involves the probabilities of all the outcomes for pairs of observables in the two sub-systems, where the outcomes are shifted by fixed amounts. The next section covers the basic matrix equation method for determining sufficiency requirements for possible Bell inequalities of the form $S\leq 4,S\leq 3,S\leq 2,S\leq 1$ in four cases for the NLHVT situation  Based on a consideration of all the sub-cases, sufficiency conditions for Bell tests in each of these four cases are set out, with the detailed derivations being presented in Appendix B. The number of sub-cases is $1,4,6,4$ for $S\leq 4,S\leq 3,S\leq 2,S\leq 1$ respectively. The conditions are in the form of inequalities involving the outcome shifts, and for the cases of $S\leq 2,S\leq 1$ , there are more than one such inequality for each of the sub-cases that are involved. At least one inequality must be satisfied for every sub-case. The last part of this section then sets out the quantum theory expressions for $S$. A comparison of the quantum theory numerical results for $S$ with the NLHVT results then follows. We again show that the CGLMP Bell inequality based on NLHVT is violated in quantum theory, reinforcing the previous conclusion based on CHSH. The paper ends with a Summary and Conclusions.
2 QTHY & HVT BASIC EQUATIONS
We consider the case of a bipartite system in which the two sub-systems may be spatialy separated. At  present we only consider one observable per sub-system- the  case of two will be considered later.
2.1 QTHY & HVT Probabilities - Bipartite Systems
The QTHY and LHVT joint probabilities $P(\alpha,\beta|\Omega_{A},\Omega_{B})$ for measurements of observables $\Omega_{A}$, $\Omega_{B}$ for sub-systems $A$, $B$ with outcomes $\alpha$, $\beta$ are:
$P(\alpha,\beta|\Omega_{A},\Omega_{B},\rho)_{Q}$ $=$ $Tr((\widehat{\Pi}_{\alpha}^{A}\otimes\widehat{\Pi}_{\beta}^{B})\widehat{\rho})$ (1)
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (2)
For QTHY $\widehat{\Pi}_{\alpha}^{A}$ is the projector onto the space of eigenvectors for operator $\widehat{\Omega}_{A}$ with eigenvalue $\alpha$. $\widehat{\rho}$ is density opr specifying the preparation process $c$ of the quantum state. For LHVT, hidden variables $\lambda$ determined with probability $P(\lambda|c)$ for preparation process $c$. The projectors satisfy $\sum_{\alpha}\widehat{\Pi}_{\alpha}^{A}=\widehat{1}^{A}$ and $\sum_{\beta}\widehat{\Pi}_{\beta}^{B}=\widehat{1}^{B}.$
In the probabilistic form of LHVT the hidden variables $\lambda$ determine each local sub-system measurement probabilities  $P(\alpha|\Omega_{A},\lambda)$, $P(\beta|\Omega_{B},\lambda)$ for classical observables $\Omega_{A}$, $\Omega_{B}$ which  are combined using classical probability theory to determine joint measurement probabilities. Since the probabilities for all outcomes must add up to unity we have $\sum_{\alpha}P(\alpha|\Omega_{A},\lambda)=1$ and $\sum_{\beta}P(\beta|\Omega_{B},\lambda)=1$. The origin of this expression for the LHVT probability expression involving the product of separate sub-system probabilities is set out below in Section 2.4. This form for the LHVT probability is widely used - apart from [5], see for example Refs. [16], [17], [9].
Both the QTHY and the LHVT probability satisfy the no-signaling condition. We have
$\sum_{\alpha}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{Q}$ $=$ $Tr((\widehat{1}^{A}\otimes\widehat{\Pi}_{\beta}^{B})\widehat{\rho})=\sum_{\alpha^{\#}}P(\alpha^{\#},\beta|\Omega_{A}^{\#},\Omega_{B},c)_{Q}$
$\sum_{\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{Q}$ $=$ $Tr((\widehat{\Pi}_{\alpha}^{A}\otimes\widehat{1}^{B})\widehat{\rho})=\sum_{\beta^{\#}}P(\alpha,\beta^{\#}|\Omega_{A},\Omega_{B}^{\#},c)_{Q}$
$\sum_{\alpha}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\beta|\Omega_{B},\lambda)=\sum_{\alpha^{\#}}P(\alpha^{\#},\beta|\Omega_{A}^{\#},\Omega_{B},c)_{LHVT}$
$\sum_{\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)=\sum_{\beta^{\#}}P(\alpha,\beta^{\#}|\Omega_{A},\Omega_{B}^{\#},c)_{LHVT}$
where two different choices of sub-system observables $\Omega_{A}^{\#},\Omega_{B}^{\#}$ and outcomes $\alpha^{\#},\beta^{\#}$ are considered. Thus the outcome for sub-system’s observable is unaffected by the outcome or choice of the other sub-system’s observable.
If the LHVT form for the probability $P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ does not apply, then what form would apply for NLHVT ? As explained in Sect. 2.4 the locality condition results in $P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ involving the product of separate sub-system outcome probabilities as in Eq (2). Hence a general NLHVT probability will be one which  does not involve separate sub-system probabilities. Thus, in the probabilistic form of this general type of NLHVT, the hidden variables $\lambda$ should determine the combined sub-system measurement probabilities  $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$, for both classical observables $\Omega_{A}$, $\Omega_{B}$  In this general non–local hidden variable theory there is no separate (or local) HVT probability for each sub-system.
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{NLHVT}=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$ (4)
Again, the sum of the probabilities for all outcomes must be unity,so $\sum_{\alpha,\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)=1$. This form for the NLHVT proability is the general form that would be involved when the local form in (2) is not applicable. Other more specific forms of NLHVT (such as Bohmian Mechanics) are briefly discussed in Sects. 1.4 and 1.5. There is no obvious reason why the no-signaling condition would apply in NLHVT.
2.2 Joint Measurement Mean Values
QTHY and LHVT expressions for the mean values of joint measurement outcomes for $\Omega_{A},\Omega_{B}$ are:
$\left\langle\Omega_{A}\otimes\Omega_{B}\right\rangle_{Q}$ $=$ $Tr(\widehat{\Omega}_{A}\otimes\widehat{\Omega}_{B})\widehat{\rho}$ (5)
$\left\langle\Omega_{A}\otimes\Omega_{B}\right\rangle_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\langle\Omega_{A}(\lambda)\right\rangle\left\langle\Omega_{B}(\lambda)\right\rangle$ (6)
where $\left\langle\Omega_{A}(\lambda)\right\rangle=\sum_{\alpha}\alpha P(\alpha|\Omega_{A},\lambda)$ is LHVT mean value for measurement of $\Omega_{A}$ when hidden variables are $\lambda$. Similarly for $\left\langle\Omega_{B}(\lambda)\right\rangle$.
The NLHVT expression for the mean values of joint measurement outcomes for $\Omega_{A},\Omega_{B}$ is
$\left\langle\Omega_{A}\otimes\Omega_{B}\right\rangle_{NLHVT}=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\langle(\Omega_{A}\otimes\Omega_{B})(\lambda)\right\rangle$ (7)
where $\left\langle(\Omega_{A}\otimes\Omega_{B})(\lambda)\right\rangle=\sum_{\alpha,\beta}\alpha\beta\,P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$ is the NLHVT mean value for measurement of $\Omega_{A}$ and $\Omega_{B}$ when hidden variables are $\lambda$.
2.3 CHSH Bell Inequality Test
The well-known CHSH Bell inequality ([6]) for bipartite system where outcomes $\alpha,\beta$ are restricted to $(+1,-1)$ and involving two measurement choices for each sub-system is
$|S|$ $=$ $|\left\langle\Omega_{A1}\otimes\Omega_{B1}\right\rangle_{LHVT}+\left\langle\Omega_{A1}\otimes\Omega_{B2}\right\rangle_{LHVT}+\left\langle\Omega_{A2}\otimes\Omega_{B1}\right\rangle_{LHVT}-\left\langle\Omega_{A2}\otimes\Omega_{B2}\right\rangle_{LHVT}\,|\,\leq 2$
Note the minus sign. This is violated for the singlet state of two spin $1/2$ systems (spin operators are in units /2Planck-constant-over-$\hbar/2$).
$\left\lfloor\Psi\right\rangle=\left(\left|+1\right\rangle_{S_{\mu}^{A}}\left|-1\right\rangle_{S_{\mu}^{B}}-\left|-1\right\rangle_{S_{\mu}^{A}}\left|+1\right\rangle_{S_{\mu}^{B}}\right)/\sqrt{2}\qquad\mu=x,y,or\;z$ (9)
for Pauli spin operators $\Omega_{A1}=\sigma_{z}^{A}$, $\Omega_{A2}=\sigma_{x}^{A}$, $\Omega_{B1}=-(\sigma_{x}^{B}+\sigma_{z}^{B})/\sqrt{2}$, $\Omega_{B2}=+(\sigma_{x}^{B}-\sigma_{z}^{B})/\sqrt{2}$, for which we find $S=2\sqrt{2}$ from quantum theory. This is based on the quantum theory mean value for the product of Pauli spin operator components along directions specified by unit vectors $\overrightarrow{a}$ and $\overrightarrow{b}$ given by $\left\langle\overrightarrow{\sigma}^{A}\cdot\overrightarrow{a}\otimes\overrightarrow{\sigma}^{B}\cdot\overrightarrow{b}\right\rangle_{Q}=-\overrightarrow{a}\cdot\overrightarrow{b}$ for the singlet state.
Note that the proof of the CHSH Bell inequality is based on the LHVT form (2) for the joint probability. It cannot be proved from the NLHVT form (4), though as we will see in Sect 3.2, the CHSH Bell inequality can be proved based on the two observable per sub-system NLHVT joint probability form (LABEL:Eq.NLHVTCGLMP).
Thus in some Bell inequalities (such as the CHSH inequalities [6]) quantum states were found in microscopic bipartite systems where failure of Bell LHVT was predicted, and experiments ([18], [19]) agreed with quantum theory. This was sufficient to rule out LHVT accounting for quantum theory in general.
2.4 Origin of LHVT Expression
The origin of the LHVT joint probability expression (2) is explained in Ref [9]. Consider two events - $A$ being the measurement of observable $\Omega_{A}$ with outcome $\alpha$ and $B$ being the measurement of observable $\Omega_{B}$ with outcome $\beta$ , where in both cases when the hidden variables are $\lambda$. These events have a joint probability $P(B,A)$ for both occuring, whilst the probability of the separate events are $P(A)$, $P(B)$. Consider a situation (such as may occur when the events occur for separate widely separated sub-systems) where  the conditional probability $P(B|A)$ for event $B$ occuring given that event $A$ occurs has the same probability as if event $A$  does not occur. This is the situation of locality. Here the conditional probability $P(B|A)$ is independent of event $A$ and thus will not depend on the probability $P(A)$ of event $A$ occuring.
$P(B|A)$ $=$ $P(B)$ (10)
$P(B,A)$ $=$ $P(B|A)P(A)$ (11)
$=$ $P(B)\times P(A)=P(A)\times P(B)=P(A,B)$ (12)
after applying Bayes’ theorem. Thus the probability $P(B,A)=P(A,B)$ of both events occuring is equal to the product of the probabilities of each separate event occuring, and we obtain the LHVT expression (2)
$P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)_{LHVT}$ $=$ $P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (13)
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (14)
after averaging over all hidden  variables $\lambda$- which occur with probability $P(\lambda|c)$. for preparation process $c$.
In the case where the hidden variable theory is non-local, the factorisation in Eq,.(13) does not apply (see Ref [16])  and $P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ would be replaced by $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$, as in Eq (4).
2.5 Quantum Entanglement and EPR Steering ?
Bell locality violation also implies that EPR steering and quantum entanglement both occur. However quantum entanglement or EPR steering can also occur for some Bell local states ([20], [16], [21], [22], [23]). Although this paper is focused on tests for Bell locality violation, Bell tests for quantum entanglement or EPR steering are also important ([23]). After all, issues regarding these effects were part of the motivation for the search for a theory to underpin quantum theory. For completeness, a brief outline of how LHVT Bell states can be divided into three categories with different featues for Quantum Entanglement and EPR Steering is set out in Appendix A. The violation of other Bell inequalities for LHVT based on this categorisation involving local hidden quantum states (LHS) [16] show whether or not Quantum entanglement or EPR steering is occuring.
Bell inequalities also exist for multi-partite systems such as in GHZ states [24] or for measurements at three different times as in  the  Leggett-Garg inequalities ([25]).  A recent review dealing with Bell correlations in macroscopic and mesoscopic systems is in Ref.[26]
In the present paper we will be focused on the so-called CGLMP Bell inequatities ([7], [8]) for bipartite systems.
3 CGLMP BELL INEQUALITIES
The CGLMP Bell inequalities will now be discussed - both for the  local hidden variable theory case and for the non-local hidden variable theory case. Here we consider the case of a bipartite system in which the two sub-systems may be spatialy separated, but now.we consider two observables per sub-system.
3.1 General CGLMP Considerations
A bipartite system with two observables per sub-system, based on a HVT joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})\equiv C(j,k,l,m)$ for all four observables $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ ([7], [8]) is considered. To shorten the notation we may write $\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}\equiv j,k,l,m$. In this bipartite case there are the same number $d$ of outcomes for each observable listed $10,1,...,d-$ 1 for each observable listed as $\ j,k,l,m$..
CGLMP ([7]) wrote their paper for a deterministic version of HVT. Here the hidden variables are the outcomes $\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}$ themselves. However,they pointed out that their theory also has a probabilistic version,- in which hidden variables $\lambda$ only determine the probabilities $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ of the outcomes. for particular $\lambda$.  In this probabilistic case we would have
$P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
For each HV choice $\lambda$ the probabilities for all outcomes must add to unity, so $\sum_{jklm}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=1$. This form for $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ applies for both the LHVT and NLHVT situations, with the (below) LHVT form (17) involving a factorisation of $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$.
There is no reason why the HVT cannot be a NLHVT whose basic probability is given by Eq (LABEL:Eq.NLHVTCGLMP) and which involves two observables per sub-system. After all, this is consistent with the idea of realism in which the hidden variables are the actual outcomes $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ of the observables $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ in the deternministic version or their probabilities $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ in the probabilistic version with  hidden variables $\lambda$. Such an approach is the basis of the.CGLMP paper. In both probabilistic and deterministic versions the outcomes have a real existence prior to any measurements, as Einstein wanted in any theory underpinning quantum theory. Also in hidden variable theory, the observables are not non-commuting Hermitian operators, so it is legitimate to have an approach where all obervables (commuting or non -commuting in quantum theory) have simultaneous outcomes. In Bohmian Mechanics the positions and momenta of every particle is ascribed a real existence. Of course for comparison with quantum theory or experiment we must construct expressions that correspond to quantitities obtainable from quantum theory formulae or which can  be measured. However this can be achieved - as for quantities such as $\left\langle\Omega_{Ai}\otimes\Omega_{Bj}\right\rangle$, which in quantum theory are given by $Tr(\widehat{\Omega}_{Ai}\otimes\widehat{\Omega}_{Bj}\;\widehat{\rho})$ or in experiment by repeated measurements on each pair of observables $\Omega_{Ai}$ and $\Omega_{Bj}$.to determine the mean value of the product of the outcomes.
CGMP also only considered cases where the outcomes for one of the observables for each sub-system was considered - $\Omega_{A1}$ or $\Omega_{A2}$ for sub-system $A$ and $\Omega_{B1}$ or $\Omega_{B2}$ for sub-system $B.$ We restrict ourselves to this situation, and hence the only HVT marginal probabilities we consider (in the general probabilistic version) are
$P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{k},\beta_{l1}|\Omega_{A2},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{j},\beta_{m}|\Omega_{A2},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{k},\beta_{m}|\Omega_{A1},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
So far we have not specified whether the CGLMP  Bell inequalities we consider are based on a local or a non-local version of HVT. [7] state that they are considering a local version of HVT. This statement is rather puzzling as CGLMP do not express $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ or $C(j,k,l,m)$ as the product of separate sub-system probabilities - which could be either as $P(\alpha_{1},\alpha_{2}|\Omega_{A1},\Omega_{A2};\lambda)P(\beta_{1},\beta_{2}|\Omega_{B1},\Omega_{B2};\lambda)$ or as $P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)$. This issue is discussed in an earlier paper on the CGLMP inequalities [27]. It was pointed out there that a theorem by Fine [28] shows that the marginal probabilities (such as $P(\alpha_{1},\beta_{1}|\Omega_{A1},\Omega_{B1})$ in Eq (LABEL:Eq.MargProb2) can be written in LHVT form. As these marginal probabilities are used to evaluate the quantity $S$ (see Eq (30) ) in the CGLMP Bell inequality, it could be claimed that CGLMP is based on  LHVT and hence a violation of the inequality shows quantum theory can not be underpinned by LHVT. That may be the case, but what we now show is that the CGLMP Bell inequality can also be established using NLHVT (just as for CHSH) - and hence the inequality - being inconsistent with quantum theory - shows that quantum theory is also not underpinned by this non-local hidden variable theory.
The joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ in the CGLMP Bell inequality (30) is applied in determining the overall result for four different sets of measurements, each involving one observable for each sub-system. These measurement choices are ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B2}$) and ($\Omega_{A1},\Omega_{B2}$).
For completeness - using the locality condition (LABEL:Eq.NLHVTCGLMP)for $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ , we see that the LHVT expression for the joint probability is given by
$P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ (17)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)$
The marginal probabilities for joint outcomes of one observable for each sub-system satisfy LHVT conditions [28]. A typical marginal probability for LHVT is $P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})$ $=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j}|\Omega_{A1},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)$ which is obviously of LHVT form. Note that $\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{m}}P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)=1$. For the NLHVT the results for the marginal probabilites are just given by (LABEL:Eq.MargProb2).
[…]
A typical CGLMP Bell inequality is
$S=P(\Omega_{A1}{\small=}\Omega_{B1})+P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}1})+P(\Omega_{A2}{\small=}\Omega_{B2})+P(\Omega_{B2}{\small=}\Omega_{A1})\leq 3$ (18)
3.2 NLHVT Proof of CHSH Bell Inequality
The proof of the CHSH Bell inequality based on local hidden variable theory is well known [6], and is set out in Ref  [9]. However, as the CHSH Bell inequality involves two observables per sub-system $\Omega_{A1},\Omega_{A2}$  for sub-system $A$ with outcomes $\alpha_{1},\alpha_{2}$ and similarly for sub-system $B,$ it is of interest to consider the situation where a general non-local hidden variable theory applies. Here there is a non-local hidden variable theory joint probability $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ for the measurement of all four observables - even though each pair of sub-system observable cannot be simultaneously measured according to quantum theory. Probability conservation gives $\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ $=1$. The question is: Does the CHSH Bell inequality still apply ?.
In this case for a probabilistic HVT
$P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ (19)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ (20)
The marginal probabilities for measurement outcomes for one observable for each sub-system would be
$P(\alpha_{1},\beta_{1}|\Omega_{A1},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{2}\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{2},\beta_{1}|\Omega_{A2},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{1}\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{1},\beta_{2}|\Omega_{A2},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{2}\beta_{1}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{2},\beta_{2}|\Omega_{A1},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{1}\beta_{1}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
where all four choices of the pairs of sub-system observables $(\Omega_{Aa},\Omega_{Bb})$ are considered. Note the sums over the unrecorded observable outcomes. These expressions are as (LABEL:Eq.MargProb2) with a notation change.
Typical terms in the expression for $S$ can then be obtained. For the first term
$\left\langle\Omega_{A1}\otimes\Omega_{B1}\right\rangle_{LHVT}$ $=$ $\sum_{\alpha_{1},\beta_{1},}\{\alpha_{1}\beta_{1}\}P(\alpha_{1},\beta_{1}|\Omega_{A1},\Omega_{B1})$
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\{\alpha_{1}\beta_{1}\}$
The other terms are derived similarly.
The quantity $S$ in the CHSH inequality would then be given by
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\{\alpha_{1}\beta_{1}+\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}-\alpha_{2}\beta_{2}\}$
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\{\alpha_{1}(\beta_{1}+\beta_{2})+\alpha_{2}(\beta_{1}-\beta_{2})\}$
In the CHSH case where $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ are all components of Pauli spin observables in spin $1/2$ sub-systems, all the outcomes are either $+1$ or $-1$. In which case $(\beta_{1}+\beta_{2})$ can either be $-2$, 00, 00 or $+2$ and for these situations $(\beta_{1}-\beta_{2})$ will be 00, $+2$, $-2$ or 00 respectively. Thus one of the two factors $(\beta_{1}+\beta_{2})$ or $(\beta_{1}-\beta_{2})$ will be zero whilst the other has a magnitude of $2$ The magnitude of both $\alpha_{1}$ and $\alpha_{2}$ is $1$ Thus in all cases |$|\{\alpha_{1}(\beta_{1}+\beta_{2})+\alpha_{2}(\beta_{1}-\beta_{2})\}|\;=2$.
Hence
$|S|\,$$\leq$ $\,\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\times 2$ (24)
$\leq$  $\,2\text{. }$
as required - using the magnitude of a sum being less than or equal to the sum of the magnitudes.
The CHSH Bell inequality can also be proved based on the LHVT form (17) for the joint probability with two observables per sub-system, where $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=P(\alpha_{1}|\Omega_{A1},\lambda)P(\alpha_{2}|\Omega_{A2},\lambda)P(\beta_{1}|\Omega_{B1},\lambda)P(\beta_{2}|\Omega_{B2},\lambda).$ So it is only for the one observable per sub-system form (4) for the NLHVT where the CHSH Bell inequality proof fails.
Thus the CHSH inequality also holds for this type of NLHVT, so its violation for the singlet state listed in Section 2.3 shows that this type of non-locality is not consistent with quantum theory. Hence we find the new result that quantum theory cannot be underpinned by this type of NLHVT. Evidently, the violation of the CHSH Bell inequality also rules out a very general type of NLHVT as well ! We will find that the same outcome applies when the CGLMP Bell inequalities are considered.
3.3 Observables and Quantum State
We now introduce the quantum theory observables that will be used in the CGLMP inequalities, along with the quantum state which will be used for comparison with HVT predictions - both local and non-local. .
The Bell inequality (18) based on the LHVT expression is violated for the quantum theory maximally entangled state
$\left|\Psi\right\rangle=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}\left|j\right\rangle_{A}\otimes\left|j\right\rangle_{B}$ (25)
for quantum theory observables whose eigenstates for eigenvalues $k$, $l$ in the case of observables $\Omega_{Aa}$ , $\Omega_{Bb}$ ($a,b=1,2$) are:
[…]
Here the $\left|j\right\rangle_{A}$ and $\left|j\right\rangle_{B}$ are orthonormal basis states for sub-systems $A$, $B$ respectively. Note that the observables $\Omega_{A1}$, $\Omega_{A2}$ are not assumed to commute, neither are $\Omega_{B1}$, $\Omega_{B2}$. These observables are alo not required to  have any obvious physical meaning. However the eigenvectors $\left|k\right\rangle_{A,a}$, $\left|l\right\rangle_{B,b}$ can be seen to satisfy the expected orthonormality conditions for different $k$, $l$. Thus
$\left\langle k\;|k^{\ast}\right\rangle_{A,a}$ $=$ $\frac{1}{d}\sum_{j=0}^{d-1}\exp i\frac{2\pi}{d}j\left(-k+k^{\ast}\right)$ (28)
$=$ $\frac{1}{d}\frac{(1-\exp i\frac{2\pi}{d}d(-k+k^{\ast}))}{(1-\exp i\frac{2\pi}{d}(-k+k^{\ast}))}=0\qquad if\text{ }k\neq k^{\ast}$
$=$ $1\qquad if\text{ }k=k^{\ast}$
with a similar result for $\left\langle l\;|l^{\ast}\right\rangle_{B,b\text{ .}}$ell non-locality for above CGLMP Bell inequalitity (18occurs for $d=2$.
The basic quantum expressions for the sub-system probabilities in the case of the maximally entangled state is (see [7])
$P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ $=$ $|(\left\langle k\;|_{A,a}\otimes\right\langle l\;|_{B,b}|)\;\left|\Psi\right\rangle|^{2}$ (29)
$=$ $\frac{1}{d}\;|\,\sum_{j=0}^{d-1}\left\langle k\;|_{A,a}|\;j\right\rangle_{A}\;|\left\langle l\;|_{B,b}|\;j\right\rangle_{B}|^{2}$
$=$ $\frac{1}{d^{3}}|\;\sum_{j=0}^{d-1}\exp-i\frac{2\pi}{d}j(k+\theta_{a})\;\exp-\frac{2\pi}{d}j(-l+\phi_{b})|^{2}$
$=$ $\frac{1}{d^{3}}\;\frac{|(1-\exp(-i2\pi(k-l+\theta_{A}+\phi_{B}))|^{2}}{|(1-\exp(-i\frac{2\pi}{d}(k-l+\theta_{A}+\phi_{B}))|^{2}}$
$=$ $\frac{1}{d^{3}}\;\frac{\sin^{2}(\pi(k-l+\theta_{a}+\phi_{b}))}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{a}+\phi_{b}))}$
$=$ $\frac{1}{2d^{3}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{a}+\phi_{b}))}$
where we have used $\sin^{2}(\pi(k-l+\theta_{a}+\phi_{b}))=\{\sin\pi(k-l)\cos\pi(\theta_{a}+\phi_{b})+(\cos\pi(k-l)\sin\pi(\theta_{a}+\phi_{b})\}^{2}=\{0+(\pm 1)(\pm 1/\sqrt{2})\}^{2}$, noting that the only $\theta_{a}+\phi_{b}$ values are $1/4,-1/4,3/4,1/4$ and $\sin\pi(1/4)=-\sin\pi(-1/4)=\sin\pi(3/4)=1/\sqrt{2}$.  Also, $(k-l)$ is always an integer.
Macroscopic Bell non-locality occurs in other CGLMP Bell inequalities for $d$ large. Experimental verification up to $d=12$ was found by [29].
3.4 NLHVT Case
We now consider a more general CGLMP Bell inequality involving
$S=P(\Omega_{A1}{\small=}\overline{\Omega_{B1}+\Delta_{11}})+P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}\Delta_{12}})+P(\Omega_{A2}{\small=}\overline{\Omega_{B2}+\Delta_{22}})+P(\Omega_{B2}{\small=}\overline{\Omega_{A1}+\Delta_{21})}$ (30)
where the outcome shifts $\Delta_{ij}$ are all integers and it is understood that $+\Delta_{ij}$ means $+\Delta_{ij}(\mathop{\mathrm{m}od}d)$.
As in the LHVT case, there are two observables $\Omega_{A1},\Omega_{A2}$ for sub-system $A$ and two observables $\Omega_{B1},\Omega_{B2}$ for sub-system $B$. The eigenvectors for all the observables are (26), (LABEL:Eq.EigenvectorsObservB)  as for the LHVT case. We will also consider quantum probabilities for the same maximally entangled state (25) that was treated in the LHVT case. However,  we now use the NLHVT expressions for the probabilities and no longer invoke locality.
First we construct the NLHVT expression for $S$. We have using the marginal probabilities in (LABEL:Eq.MargProb2)
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})\,\delta_{j,l+\Delta_{11}}+\mathop{\displaystyle\sum}\limits_{k,l}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1})\,\delta_{l,k+\Delta_{12}}$ (35)
$+\mathop{\displaystyle\sum}\limits_{k,m}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2})\,\delta_{k,m+\Delta_{22}}+\mathop{\displaystyle\sum}\limits_{j,m}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2})\,\delta_{m,j+\Delta_{21}}$
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,l}\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\,\delta_{j,l+\Delta_{11}}\\ +\mathop{\displaystyle\sum}\limits_{k,l}\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\,\delta_{l,k+\Delta_{12}}\\ +\mathop{\displaystyle\sum}\limits_{k,m}\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\,\delta_{k,m+\Delta_{22}}\\ +\mathop{\displaystyle\sum}\limits_{j,m}\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\,\delta_{m,,j+\Delta_{21}}\end{array}\right\}$
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{jklm}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\\ \times\{\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}}\}\end{array}\right\}$ (38)
Clearly the quantity $\delta=\{\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}}\}$ can have values $4,3,2,1$ depending on how many of the Kronecka deltas are equal to $1$ rather than  00. We see that the CGLMP Bell inequality in  the NLHVT case will just depend on the value for $\delta$
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{jklm}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\;\times\delta\right\}$ (39)
$\leq$ $\delta$
using $\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=1$ and $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)=1$. In the next Section we determine sufficiency requirements on the shifts $\Delta_{ij}$ for the cases $\delta=4,3,2,1$.
By comparison with (38) the correspondiing expression if LHVT applies is
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)\\ \times(\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}})\end{array}\right\}$ (42)
where here the locality based factorisation of $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ occurs. However the same sum of Kronecka deltas is present. In this LHVT version we see that we also have
$S\leq\delta$ (44)
using $\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)=1$ as each factor is equal to $1$, for example $\sum_{j}P(\alpha_{j}|\Omega_{A1},\lambda)=1$.
This means that if the NLHVT is violated by quantum theory, then the LHVT will also be violated.
Note that the sum of the joint measurement probabilities where specific relationships occur between the measurement outcomes for the two sub-systems is in general less that the sum of the joint measurement probabilities for all such relationships. Thus $P(\Omega_{A1}{\small=}\overline{\Omega_{B1}+\Delta_{11}})=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,\delta_{j,l+\Delta_{11}}\leq\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{j,l}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)=P(\Omega_{A1},\Omega_{B1})=1$., where $P(\Omega_{A1},\Omega_{B1})$ is the overall probability for all outcomes of joint measurements for the choice $\Omega_{A1}$ and $\Omega_{B1}$. Similar considerations apply to $P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}\Delta_{12}})$, $P(\Omega_{A2}{\small=}\overline{\Omega_{B2}+\Delta_{22}})$ and $P(\Omega_{B2}{\small=}\overline{\Omega_{A1}+\Delta_{21})\text{.}}$.
4 REQUIREMENTS FOR NLHVT BELL INEQUALITIES
First we consider the conditions involving the Kroneka deltas involved in the simple CGLMP inequality and the matrix approach used to treat them. Since the requirements only involve considering the situations where the Kronecka sum $\delta$ is given by $4,3,2$ or $1$, it follows that the requirements will be the same for both the LHVT and NLHVT cases.
4.1 Conditions and Basic Matrix Approach
To derive the sufficiency requirements on the outcome shifts $\Delta_{ij}$ for the NLHVT Bell Inequalities (39) we must consider the conditions where the various Kronecka deltas are equal to unity. These conditions are
$C1$ ::: $\qquad j=l+\Delta_{11}$
$C2$ ::: $\qquad l=k+\Delta_{12}$
$C3$ ::: $\qquad k=m+\Delta_{22}$
$C4$ ::: $\qquad m=j+\Delta_{21}$ (45)
We will consider cases where; Case 1 - all of $C1,C2,C3,C4$ are true (T); Case 2 - three of $C1,C2,C3,C4$ are true and one is false (F); Case 3 - two of $C1,C2,C3,C4$ are true and two are false; Case 4 - one of $C1,C2,C3,C4$ is true and three are false. For these four Cases, there are a number of Sub-Cases depending on which of $C1,C2,C3,C4$ are true and which are false - one for Case 1, four for Case 2, six for Case 3 and four for Case 4. For the four Cases the value of $\delta=$ $(\delta_{j,l+\Delta_{11}}+\delta_{l,k+\Delta_{12}}+\delta_{k,m+\Delta_{22}}+\delta_{m,j+\Delta_{21}})$ are the same for every Sub-Case, namely $\delta=4,3,2,1$. for Cases $1,2,3,4$ respectively. As described in Sect.3.4 for these four Cases the values of $\delta$ are used to determine the Bell inequality $S\leq\delta$ (39) using Eq. (38) for $S$ (defined in Eq. (30)). However, the requirements on the shifts $\Delta_{ij}$ for each Sub-Case will be different, and these are set out below. Details are given in Appendix B. .
To derive these requirements it turns out that we must consider matrix equations of the form
$M\;\mathbf{v=\Delta}$ (46)
where $M$ is a square matrix and $\mathbf{v}$ is a column vector whose elements are some (or all) of the integers $j,k,l,m$ and $\mathbf{\Delta}$ is a column vector whose elements are sums of the various $\Delta_{ij}$. We will use the result that if the determinent of $M$ is zero ($|M|=0$) then there is only a possible solution for $\mathbf{v}$ if all the elements of the corresponding $\mathbf{\Delta}$ are zero. This will then lead to a set of requirements on the shifts $\Delta_{ij\text{ }}$to lead to Bell inequalities: $S\leq 4$, $S\leq 3$, $S\leq 2$, $S\leq 1$ for each Sub-Case of the Cases 1, 2, 3, 4. Note that we merely have to have one element of $\Delta$ to be non-zero to show there is no solution for $\mathbf{v}$.
The key point is that the CGLMP Bell Inequalities are of the form $S\leq\delta$, so they only depend on the number of the conditions $C1,C2,C3,C4$ that are satisfied. The basic strategy for each Sub Case is to take each condition that is required to be False, then substitute the conditions that are required to be True into these to arrive at the matrix equation (46) for the remaining indices from $j,k,l,m$ and then identify the column vector $\Delta$ for this  Sub-Case. We then show that the determinent $|M|$ is zero. As then there is to be no solution for the conditions that are False, we then obtain the requirement that at least one of  the elements of $\Delta$ is non-zero. Hence the requirement for each Sub-Case will be a list of And/Or statements. Each statement is a set of And/Or inequalities (or an equality in the Case of $S\leq 4$) arranged in rows for the various Sub-Cases, and involving the outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$. If at least one of these these inequalities are satisfied in every row, then we have obtained sufficiency requirements to show $S\leq\delta$ for the Case involved. Note that we do not obtain necessity requirements, as a particular set of outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ may provide sufficiency requirements for more than one of $S\leq 3$, $S\leq 2$ and  $S\leq 1$.  If more than one is satisfied then the overall NLHVT prediction is $S$leq the smallest of $3,2,1$ satisfied. This is logical since if for the outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ we can show that $S\leq 3$, $S\leq 2$ and $S\leq 1$ then overall we have $S\leq 1$.  The Case of $S\leq 4$ is an  exception, as we will see. Sufficiency requirements for $S\leq 4$ are incompatible with those for $S\leq 3$, $S\leq 2$ and  $S\leq 1$.
4.2 Requirements for Bell Inequality Tests
These Sub-Case requirements are derived in Appendix B for all Sub-Cases of the four Cases. These are summarised below in Eqs. (47), (48) , (LABEL:Eq.RequireLessTwoSubCases) and (LABEL:Eq.RequireLessOneSubCases). The resulting requirements for the four Cases is based on the distinct Sub-Case requirements and must cover all Sub-Cases to constitute sufficient (rather than necessary) requirements for concluding that according to our general NLHVT we have $S$ satisfying $S\leq 4$, $S\leq 3$, $S\leq 2$ or $S\leq 1$. These requirements are set out below in Eqs. (47), (48), (LABEL:Eq.RequireLessTwoSubCases) and (LABEL:Eq.RequireLessOneSubCases) respectively. In deriving these results we have used $\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=1$ and $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)=1$.
4.2.1 Case1 - S4absent4$\leq$ 4≤ 4
The requirements for the single Sub-Case are
$\Delta_{11}=0\quad And\quad\Delta_{12}=0\quad And\quad\Delta_{22}=0\quad And\quad\Delta_{21}=0$ (47)
If all these requirements are satisfied the Bell inequality $S\leq 4$ would be guaranteed. Note that if these requirements are satisfied then those for $S\leq 3,S\leq 2,S\leq 1$ are not.
4.2.2 Case 2 - S $\leq 3$
The requirements for all four Sub-Cases are the same
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0\qquad Sub-Cases\text{ }a,b,c,d$ (48)
If this requirement applies the Bell inequality $S\leq 3$ would be satisfied.
4.2.3 Case 3 - S $\leq 2$
The requirements for the six different Sub-Cases are
$\Delta_{22}$$\neq$ $0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{22}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }b$
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{22}+\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0\qquad\qquad Sub-Case\text{ }e$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0\qquad Sub-Case\text{ }f$
Since we wish to guarantee that the Bell inequality $S\leq 2$ would occur irrespective of the Sub-Case, we can be certain of this by checking for a given set of outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ that at least one of the distinct inequalities in each row of the last equation is satisfied.  For exanple, the requirements $\Delta_{11}\neq 0,\Delta_{12}\neq 0,\Delta_{22}\neq 0$ and $\Delta_{21}\neq 0$ would satisfy the requirements for Sub-Cases $(a),(c),(d)$ and $(f)$ but not those for Sub-Cases $(b)$ and $(e)$, so with these four requirements alone we cannot conclude that $S\leq 2$.
4.2.4 Case 4 - S $\leq 1$
The requirements for the four different Sub-Cases are
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{11}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }b$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{12}\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
Since we wish to guarantee that the Bell inequality $S\leq 1$ would occur irrespective of the sub-case, we can be certain of this by checking for a given set of outcome shifts $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ that at least one of  the distinct inequalities in each row of the last equation are satisfied.
4.3 Example
As an example of what happens for a particular choice of outcome shifts we choose the case $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(1,-2,1,-2).$ It is obvious that the requirement for $S\leq 4$ is not satisfied.
We substitute the numerical values into Eqs. (48), (LABEL:Eq.RequireLessTwoSubCases) and (LABEL:Eq.RequireLessOneSubCases) .
For $S\leq 3$ (all Sub-Case) we see that $\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}=(-2)\neq 0$ as required.
For $S\leq 2$we have
$\Delta_{22}$ $=$ $(2)\neq 0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}=(-3)\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{22}+\Delta_{12}$ $=$ $(-1)\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}=(-1)\neq 0\qquad\qquad Sub-Case\text{ }b$
$\Delta_{12}$$\neq$ $(-2)\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}=(0)\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{22}+\Delta_{12}+\Delta_{11}$ $=$ $(0)\neq 0\quad And/Or\quad\Delta_{21}=(-2)\neq 0\qquad\qquad Sub-Case\text{ }d$
$\Delta_{12}+\Delta_{11}$ $=$ $(-1)\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}=(-1)\neq 0\qquad\qquad Sub-Case\text{ }e$
$\Delta_{11}$ $=$ $(1)\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}=(-3)\neq 0\qquad Sub-Case\text{ }f$
so as there is at least one  inequality satisfied for each Sub-Case we see that the requirements for $S\leq 2$are satisfied.
For $S\leq 1$we have
$\Delta_{12}$ $=$ $(-2)\neq 0\quad And/Or\quad\Delta_{22}=(1)\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}=(-1)\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{11}+\Delta_{12}$ $=$ $(-1)\neq 0\quad And/Or\quad\Delta_{22}=(1)\neq 0\quad And/Or\quad\Delta_{21}=(-2)\neq 0\qquad Sub-Case\text{ }b$
$\Delta_{11}$ $=$ $(1)\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{12}=(-1)\neq 0\quad And/Or\quad\Delta_{21}=(-2)\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{11}$ $=$ $(1)\neq 0\quad And/Or\quad\Delta_{12}=(-2)\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{21}=(-1)\neq 0\qquad\qquad Sub-Case\text{ }d$
so as there is at least one  inequality satisfied for each Sub-Case we see that the requirements for $S\leq 1$are satisfied.
Thus, the requirements for $S\leq 3$, $S\leq 2$ and $S\leq 1$.are all satisfied and we can conclude overall that NLHVT predicts $S\leq 1$ for this choice of outcome shifts.
4.4 Quantum Theory Test
Here we compare the quantum theory expressions for $S$ with those from NLHVT based on a specic choice for the $\Delta_{ij}$.
In QTHY the result for the quantity $S$ is given by
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{jl}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\rho)_{Q}\;\delta_{j,l+\Delta_{11}}+\mathop{\displaystyle\sum}\limits_{kl}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\rho)_{Q}\;\delta_{l,k+\Delta_{12}}$
$+\mathop{\displaystyle\sum}\limits_{km}P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\rho)_{Q}\;\delta_{k,m+\Delta_{22}}+\mathop{\displaystyle\sum}\limits_{jm}P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\rho)_{Q}\;\delta_{m,j+\Delta_{21}}$
and using the QTHY result (29) for $P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ in the case of the maximally entangled state we obtain
$S$ $=$ $\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jl}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-l+\theta_{1}+\phi_{1}))}\;\delta_{j,l+\Delta_{11}}$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{kl}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{2}+\phi_{1}))}\;\delta_{l,k+\Delta_{12}}$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{km}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-m+\theta_{2}+\phi_{2}))}\;\delta_{k,m+\Delta_{22}}$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jm}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-m+\theta_{1}+\phi_{2}))}\;\delta_{m,j+\Delta_{21}}$
$=$ $\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{l}\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{11}+\theta_{1}+\phi_{1}))}\;+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{l}\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{12}+\theta_{2}+\phi_{1}))}\;$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{m}\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{22}+\theta_{2}+\phi_{2}))}\;+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{m}\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{21}+\theta_{1}+\phi_{2}))}\;$
$=$ $\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{11}+\theta_{1}+\phi_{1}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{12}+\theta_{2}+\phi_{1}))}$
$+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\Delta_{22}+\theta_{2}+\phi_{2}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-\Delta_{21}+\theta_{1}+\phi_{2}))}$
for $S$ in the QTHY approach. Note that adding an integer multiple of $d$ to the $\Delta_{ij}$ does not alter the result.
Note that although the four sets of measurements for the result for $S$ involve one pair of the two sub-system observables at a time - ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A1},\Omega_{B1}$) or  ($\Omega_{A1},\Omega_{B1}$) the outcomes for the other pair of observables are disregarded. In an earlier paper [8] we pointed out that although for HVT (which involves non-quantum observables) the way the other pair of observable outcomes was disregared was unambiguous, this was not the case in quantum theory. Options included (a) not measuring the other pair  observables at all (b) measuring the other pair of observables and then discarding the results - and this could be done in two ways depending on whether the disregarded observables were measured first or second. The quantum expressions differ for (a) and (b) since the latter involve a projection operation into the space of the disregarded observables. Any of these would be a possible quantum prediction for comparison purposes. However, as it is only necessary to show HVT is inconsistent with one possible quantum prediction we choose the simplest option, namely that in each term the other observables are not measured at all.  This was the choice made by CGLMP and Eq (29) for the quantum theory expression $P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ is determined this way.
We could choose as a special case $\Delta_{ij}$ in the form $\Delta_{ij}=d\times\Delta_{ij}^{\ast}$ where the $\Delta_{ij}^{\ast}$ are all integers. In this case, using $\sin^{2}(\pi\Delta_{ij}^{\ast}+\eta)=(\sin(\pi\Delta_{ij}^{\ast})\cos(\eta)+\cos(\pi\Delta_{ij}^{\ast})\sin(\eta))^{2}=\sin^{2}(\eta)$ in Eq (LABEL:Eq.QThyResultS_CGLMP), we obtain for the QTHY approach
$S$ $=$ $\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(d\Delta_{11}^{\ast}+\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-d\Delta_{12}^{\ast}+\frac{3}{4}))}$ (55)
$+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(d\Delta_{22}^{\ast}+\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-d\Delta_{21}^{\ast}-\frac{1}{4}))}$
$=$ $\frac{1}{2d^{2}}\;\left(\frac{1}{\sin^{2}(\frac{\pi}{4d})}+\frac{1}{\sin^{2}(\frac{3\pi}{4d})}+\frac{1}{\sin^{2}(\frac{\pi}{4d})}+\frac{1}{\sin^{2}(\frac{\pi}{4d})}\right)$
$=$ $\frac{1}{2d^{2}}\;\left(\frac{3}{\sin^{2}(\frac{\pi}{4d})}+\frac{1}{\sin^{2}(\frac{3\pi}{4d})}\right)$
We have used the values $a=1,2\quad\theta_{1}=0,\theta_{2}=1/2$ and $b=1,2\quad\phi_{1}=1/4,\phi_{2}=-1/4$ from Eqs (26), (LABEL:Eq.EigenvectorsObservB) for  the eigenvectors. This last result shows that in quantum theory the value of $S$ is independent of the $\Delta_{ij}^{\ast}$. It only depends on the integer $d$. The NLHVT Bell inequalities for $S$ do of course depend on the $\Delta_{ij}^{\ast}$.
For the choice $d=10$ we find that in QTHY $S\approx 2.52$, so this exceeds $S=2$ and $S=1$ though not $S=4$ and $S=3$ that could occur in NLHVT depending on the requirements for these cases being met. For the choice $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-2,1,-2)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3$, for $S\leq 2$ and for $S\leq 1$ are all met. . Other choices are also be made. For example with $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-3,1,-3)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3,S\leq 2$ and $S\leq 1$ are all met. In  both these  situations NLHVT predicts $S\leq 1$ this being the smallest of the three inequalities. Thus we have found cases for $d=10$ where the quantum theory prediction that $S=2.52$ is inconsistent with the NLHVT Bell inequality prediction of $S\leq 1.$
We could also choose $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ as in the original CGLMP Bell inequality (18). For this choice the requirements for $S\leq 3$, for $S\leq 2$ and for $S\leq 1$ are all met. .In this case the CGLMP Bell inequality predicts $S\leq 1$. For this choice the QTHY expression for $S$ is given by
$S$ $=$ $\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-1+\frac{3}{4}))}$ (56)
$+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(\frac{1}{4}))}\;+\frac{1}{2d^{2}}\;\frac{1}{\sin^{2}(\frac{\pi}{d}(-\frac{1}{4}))}$
$=$ $\frac{2}{d^{2}}\frac{1}{\sin^{2}(\frac{\pi}{4d})}$
For the case of $d=2$ this gives $S=3.41$. For the choice $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ the requirements for $S\leq 4$ are not met, but those for $S\leq 1$ are met,.However NLHVT is inconsistent with QTHY in this choice of $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ as it was for LHVT where in both  types of HVT the inequality $S\leq 1$ is violated in Q THY.
5 SUMMARY AND CONCLUSION
A brief review of the origins of local hidden variable theory, Bell local states and Bell inequalities has been presented. We have also outlined the basic justification for LHVT. The classification of Bell local states into three categorise based on the local hidden quantum state approach of Wiseman [16] has been explained. We outlined how the Bell inequalities provide tests for whether local hidden variable theory is in conflict with quantum theory and also which categories of Bell local states exhbit quantum entanglement or EPR steering.
The new result in this paper is that a Bell-type non-local hidden variable theory has been presented, and Bell inequalities obtained based on this NLHVT. This is based on the approach of Collins, Gisin, Linden, Massar Popescu (CGLMP) for bipartite systems, where the basic joint probability involves the outcomes for two observables for each  sub-system  In addition to showing that the well-known CHSH inequality can also be proved  for such a NLHVT, we have also considered CGLMP Bell inequalities involving the quantity $S$ that involves joint probabilities for outcomes from a pair of sub-system observables,  one in each sub-system. Sufficiency conditions for the outcome shifts have been obtained for CGLMP inequalities $S\leq 4,S\leq 3,S\leq 2$ and $S\leq 1$. It has been shown that in a maximally entangled state in a bipartite system clear quantum theory violations of these CGLMP Bell inequalities occur , thus showing that it is not posible to underpin or complete quantum theory via a CGLMP  Bell-type non-local hidden variable theory. The same sufficiency conditions also show that there is also a quantum theory violation for the CGLMP version of LHVT.
It should be pointed out that the observables chosen have no obvious physical meaning. Also the experimental measurements for the NLHVT Bell inequality terms would difficult to perform. Nor would the maximally entangled state be easy to prepare. However these issues are irrelevant as all that is needed is that the observables are possible quantum observables, the state is a possible quantum state and the Bell inequality measurements could be carried out.
The result that a Bell-type non-local hidden variable theory is also in conflict with quantum theory is of some significance since it shows that quantum theory can not be understood  in terms of Bell-type hidden variable theory. This is irrespective of whether the hidden variable theory is local or non-local.
6 ACKNOWLEDGEMENTS
The author wishes to acknowledge discussions on quantun foundations with several colleagues - M Reid, S Barnett, B Garraway, L Heaney, J Goold, T Busch, L Rosales, R Teh, B Opanchuk and P Drummond.
7 APPENDIX A - LOCAL HIDDEN QUANTUM STATES, ENTANGLEMENT & EPR STEERING
Quantum states can be divided into entangled and separable states, but in LHVT can also be divided into Bell local and Bell non-local states. These categories overlap [20].
7.1 Separable States
For separable states, prepare sub-systems $A$, $B$ in quantum states $\widehat{\rho}_{R}^{A}$, $\widehat{\rho}_{R}^{B}$, where combined state $\widehat{\rho}_{R}^{A}\otimes\widehat{\rho}_{R}^{B}$ occurs with classical probability $P_{R}$ [20]. With density opr $\widehat{\rho}_{sep}=\sum_{R}P_{R}\;\widehat{\rho}_{R}^{A}\otimes\widehat{\rho}_{R}^{B}$ and $P(\alpha|\Omega_{A},R)=Tr_{A}(\widehat{\Pi}_{\alpha}^{A}\widehat{\rho}_{R}^{A})$ etc.,joint outcome prob is given by a LHVT form:
$P(\alpha,\beta|\Omega_{A},\Omega_{B},\rho_{sep})_{Q}=\sum_{R}P_{R}\,P(\alpha|\Omega_{A},R)\times P(\beta|\Omega_{B},R)$ (57)
This is of the same form as for Bell local states, showing that all separable states are Bell-local ($R\rightarrow\lambda$). Hence Bell non-local states must be entangled.
However, some Bell local states are also entangled  [20]. For Bell-local states which are separable, both sub-systems are associated with a quantum density operator. Hence idea of a local hidden quantum state (LHQS) was introduced [16], [21], [22].
7.2 Local Hidden Quantum State
Bell-local states in bipartite systems can be further divided into three disjoint categories, depending on two, one or none of the sub-systems being associated with a local hidden quantum state [16], [21], [22]. LHQS $\widehat{\rho}_{C}(\lambda)$ are to be possible quantum states -- these must comply with conditions such as the super-selection rules of there being no coherences between states with differing (massive) particle numbers.
7.3 Bell-Local States - Three Categories
For the bipartite case - there are three categories of Bell-local states
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\textstyle\sum}\limits_{\lambda}P_{Q}(\alpha|\Omega_{A},\lambda)\,P_{Q}(\beta|\Omega_{B},\lambda)\,P(\lambda|c)\qquad Cat\;1$
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\textstyle\sum}\limits_{\lambda}P(\alpha|\Omega_{A},\lambda)\,P_{Q}(\beta|\Omega_{B},\lambda)\,P(\lambda|c)\qquad Cat\;2$
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\textstyle\sum}\limits_{\lambda}P(\alpha|\Omega_{A},\lambda)\,P(\beta|\Omega_{B},\lambda)\,P(\lambda|c)\qquad Cat\;3$
where subscript $Q$ indicates sub-system measurement probability determined from a local hidden quantum state $\widehat{\rho}_{C}(\lambda)$, via $P_{Q}(\gamma|\Omega_{C},\lambda)=Tr_{C}(\widehat{\Pi}_{\gamma}^{C}\widehat{\rho}_{C}(\lambda))$. Bell inequalities will differ depending on whether there are sub-systems associated with a LHQS.
7.4 Tests for Quantum Entanglement and EPR Steering
For the case of two modes. we introduce spin operators $S_{x}=\frac{1}{2}(b^{{\dagger}}a+a^{{\dagger}}b),S_{y}=\frac{1}{2i}(b^{{\dagger}}a-a^{{\dagger}}b),S_{z}=\frac{1}{2}(b^{{\dagger}}b-a^{{\dagger}}a)$.
Category 1 states are same as separable states since both sub-system probabilities are given by quantum expressions. Violation of inequalities based on Category 1 states $\rightarrow$ Quantum entanglement (non-separability) occurs. Example: Spin squeezing test for two mode entanglement [30] $\left\langle\Delta S_{x}^{2}\right\rangle<\frac{1}{2}|\left\langle S_{z}\right\rangle$ or $|\left\langle\Delta S_{y}^{2}\right\rangle<\frac{1}{2}|\left\langle S_{z}\right\rangle|$.
For Category 2 states only one sub-system probability ($B$) is given by quantum expression involving LHS quantum density operator -- $\widehat{\rho}^{B}(\lambda)$. Here no EPR steering of $B$ by $A$ occurs $\rightarrow$ Outcomes $\beta$ for $\Omega_{B}$ depend only on $\widehat{\rho}_{B}(\lambda)$- not on $\alpha,\Omega_{A}$. Violation of inequalities based on Category 2 states $\rightarrow$ EPR steering of $B$ outcomes by $A$ measurement occurs. Example: generalised Hillary-Zubairy planar spin variance test for two mode EPR steering [31], [23]. $\left\langle\Delta S_{x}^{2}\right\rangle+\left\langle\Delta S_{y}^{2}\right\rangle<(\frac{1}{4}\left\langle N\right\rangle-\frac{1}{2}\left\langle S_{z}\right\rangle)$.
7.5 Tests for Bell Non-Locality
For Category 3 states neither of sub-system probabilities given by quantum expression - no local hidden quantum state. . Inequality violations based on Category 3 states $\rightarrow$ Bell locality cannot apply. Example: CHSH Bell Inequality (above) for two spin 1/2 systems. Thus Bell non-locality is shown by violations of inequalities based on
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (59)
7.6 Bell - Non-Local States
A fourth category of states are the Bell non-local states.
$P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{NLHVT}=\mathop{\textstyle\sum}\limits_{\lambda}P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)\,\,P(\lambda|c)\qquad Cat\;4$ (60)
Here there is no separate $P(\gamma|\Omega_{C},\lambda)\,$for each sub-system. Violation of inequalities based on Category 4 states show that non-local hidden variable theories also fail.
7.7 Overall Scheme
The overall scheme can be pictured as (see [27], [23])  Here we have assumed that Quantum Theory is underpinned by Hidden Variable Theory - local or non-local and that Bell non-local states are quantum entangled states. The issue of whether or not quantum theory is underpinned by a non-local hidden variable theory is the focus of the main part of this paper.
Figure 1: The classifcation of hidden variable Bell local states into three categories, plus  the hidden variable Bell non-local states. Also the classification of quantum states as separable and entangled. The features of the various types of states are stated.
8 APPENDIX B - REQUIREMENTS FOR BELL TESTS
In this Appendix  we will derive the requirements for Bell Inequality tests in the NLHVT situation.
8.1 Case 1 - Bell Inequality - All Conditions T
If $C1=T,C2=T,C3=T,C4=T$ we then have
$\left[\begin{array}[]{cccc}1&0&-1&0\\ 0&-1&1&0\\ 0&1&0&-1\\ -1&0&0&1\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\\ l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{12}\\ \Delta_{22}\\ \Delta_{21}\end{array}\right]$ (61)
As $|M|=0$ there is only a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}=0\quad\Delta_{12}=0\quad\Delta_{22}=0\quad\Delta_{21}=0$ (62)
This solution is $j=k=l=m$ and here we have $\delta=4$
Hence applying this result for $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(4\right)\,\right\}$ (63)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(4)$
$\leq$ 44 44
The Bell inequality is then $S\leq 4$.
8.2 Case 2 - Bell Inequality - Three Conditions T, One F
Here there are four sub-cases. In all of thes sub-cases we have $\delta=3.$
8.2.1 Sub-Case a - $C1=T,C2=T,C3=T,C4=F$
Substituting $C1,C2,C3$ into $C4$ we have
$m$ $=$ $j+\Delta_{21}$ (64)
$=$ $l+\Delta_{11}+\Delta_{21}$
$=$ $k+\Delta_{12}+\Delta_{11}+\Delta_{21}$
$=$ $m+\Delta_{22}+\Delta_{12}+\Delta_{11}+\Delta_{21}$
As $C4=F$ it follows that the last equation is false. Hence the following requirement must be satisfied
$\Delta_{22}+\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0$ (65)
8.2.2 Sub-Case b - $C1=T,C2=T,C3=F,C4=T$
Substituting $C1,C2,C4$ into $C3$ we have
$k$ $=$ $m+\Delta_{22}$ (66)
$=$ $j+\Delta_{21}+\Delta_{22}$
$=$ $l+\Delta_{11}+\Delta_{21}+\Delta_{22}$
$=$ $k+\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}$
As $C3=F$ it follows that the last equation is false. Hence the following requirement must be satisfied
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (67)
8.2.3 Sub-Case c - $C1=T,C2=F,C3=T,C4=T$
The requirement and the Bell inequality is the same as for Sub-Cases a and b.
8.2.4 Sub-Case d - $C1=F,C2=T,C3=T,C4=T$
The requirement and the Bell inequality is the same as for Sub-Cases a, b and c.
8.2.5 All Sub-Cases of Case 2
Overall, we see that if the requirement is satisfied, we can say $\delta=3.$
Hence applying this result for $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(3\right)\,\right\}$ (68)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(3)$
$\leq$ 33 33
From above the following requirement must be satisfied for Case 2 to cover all sub-cases.
$\Delta_{12}+\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (69)
The Bell inequality is then $S\leq 3$.
8.3 Case 3 - Bell Inequality - Two Conditions T, Two Conditions F
Here there are six sub-cases. In all of these sub-cases $\delta=2$.
8.3.1 Sub-Case a $C1=T,C2=T,C3=F,C4=F$
Substituting $C1,C2$ into $C3,$ $C4$ we have
$k$ $=$ $m+\Delta_{22}$
$m$ $=$ $j+\Delta_{21}$ (70)
$=$ $l+\Delta_{11}+\Delta_{21}$
$=$ $k+\Delta_{12}+\Delta_{11}+\Delta_{21}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}k\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{22}\\ \Delta_{12}+\Delta_{11}+\Delta_{21}\end{array}\right]$ (71)
Since $C3=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{22}\neq 0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0$ (72)
8.3.2 Sub-Case b $C1=T,C2=F,C3=T,C4=F$
Substituting $C1,C3$ into $C2,$ $C4$ we have
$l$ $=$ $k+\Delta_{12}$
$=$ $m+\Delta_{22}+\Delta_{12}$
$m$ $=$ $j+\Delta_{21}$ (73)
$=$ $l+\Delta_{11}+\Delta_{21}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{22}+\Delta_{12}\\ \Delta_{11}+\Delta_{21}\end{array}\right]$ (74)
Since $C2=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0$ (75)
8.3.3 Sub-Case c $C1=T,C2=F,C3=F,C4=T$
Substituting $C1,C4$ into $C2,$ $C3$ we have
$l$ $=$ $k+\Delta_{12}$
$k$ $=$ $m+\Delta_{22}$ (76)
$=$ $j+\Delta_{21}+\Delta_{22}$
$=$ $l+\Delta_{11}+\Delta_{21}+\Delta_{22}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}-1&+1\\ +1&-1\end{array}\right]\;\left[\begin{array}[]{c}k\\ l\end{array}\right]=\left[\begin{array}[]{c}\Delta_{12}\\ \Delta_{11}+\Delta_{21}+\Delta_{22}\end{array}\right]$ (77)
Since $C2=F,C3=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{12}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0$ (78)
8.3.4 Sub-Case d $C1=F,C2=T,C3=T,C4=F$
Substituting $C2,C3$ into $C1,$ $C4$ we have
$j$ $=$ $l+\Delta_{11}$
$=$ $k+\Delta_{12}+\Delta_{11}$
$=$ $m+\Delta_{22}+\Delta_{12}+\Delta_{11}$
$m$ $=$ $j+\Delta_{21}$ (79)
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{22}+\Delta_{12}+\Delta_{11}\\ \Delta_{21}\end{array}\right]$ (80)
Since $C1=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{22}+\Delta_{12}+\Delta_{11}\neq 0\quad And/Or\quad\Delta_{21}\neq 0$ (81)
8.3.5 Sub-Case e $C1=F,C2=T,C3=F,C4=T$
Substituting $C2,C4$ into $C1,$ $C3$ we have
$j$ $=$ $l+\Delta_{11}$
$=$ $k+\Delta_{12}+\Delta_{11}$
$k$ $=$ $m+\Delta_{22}$ (82)
$=$ $j+\Delta_{21}+\Delta_{22}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}+1&-1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\end{array}\right]=\left[\begin{array}[]{c}\Delta_{12}+\Delta_{11}\\ \Delta_{21}+\Delta_{22}\end{array}\right]$ (83)
Since $C1=F,C3=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{12}+\Delta_{11}\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0$ (84)
8.3.6 Sub-Case f $C1=F,C2=F,C3=T,C4=T$
Substituting $C3,C4$ into $C1,$ $C2$ we have
$j$ $=$ $l+\Delta_{11}$
$l$ $=$ $k+\Delta_{12}$ (85)
$=$ $m+\Delta_{22}+\Delta_{12}$
$=$ $j+\Delta_{21}+\Delta_{22}+\Delta_{12}$
Thus we have the marix equation
$\left[\begin{array}[]{cc}-1&+1\\ -1&+1\end{array}\right]\;\left[\begin{array}[]{c}l\\ j\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{21}+\Delta_{22}+\Delta_{12}\end{array}\right]$ (86)
Since $C1=F,C2=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0$ (87)
8.3.7 All Sub-Cases of Case 3
Overall, we see that if the requirement is satisfied, we can say $\delta=2$.
Hence applying this value of $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(2\right)\,\right\}$ (88)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(2)$
$\leq$ 22 22
From above the following requirements must be satisfied for Case 3 to cover all sub-cases.
$\Delta_{22}$$\neq$ $0\quad And/Or\quad\Delta_{12}+\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{22}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }b$
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{11}+\Delta_{21}+\Delta_{22}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{22}+\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
$\Delta_{12}+\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0\qquad\qquad Sub-Case\text{ }e$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{21}+\Delta_{22}+\Delta_{12}\neq 0\qquad Sub-Case\text{ }f$
The Bell inequality is then $S\leq 2$.
8.4 Case 4 Bell Inequality - One Condition T, Three Conditions F
Here there are four sub-cases. In all these  sub-cases  $\delta=1$.
8.4.1 Sub-Case a $C1=T,C2=F,C3=F,C4=F$
Substituting $C1$ into $C2,$ $C3,$ $C4$ we have
$l$ $=$ $k+\Delta_{!2}$
$k$ $=$ $m+\Delta_{22}$
$m$ $=$ $j+\Delta_{21}$ (90)
$=$ $l+\Delta_{11}+\Delta_{21}$
Thus we have the marix equation
$\left[\begin{array}[]{ccc}-1&+1&0\\ +1&0&-1\\ 0&-1&+1\end{array}\right]\;\left[\begin{array}[]{c}k\\ l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{!2}\\ \Delta_{22}\\ \Delta_{11}+\Delta_{21}\end{array}\right]$ (91)
Since $C2=F,C3=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{!2}\neq 0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0$ (92)
8.4.2 Sub-Case b $C1=F,C2=T,C3=F,C4=F$
Substituting $C2$ into $C1,$ $C3,$ $C4$ we have
$j$ $=$ $l+\Delta_{11}$
$=$ $k+\Delta_{!2}+\Delta_{11}$
$k$ $=$ $m+\Delta_{22}$
$m$ $=$ $j+\Delta_{21}$ (93)
Thus we have the marix equation
$\left[\begin{array}[]{ccc}+1&-1&0\\ 0&+1&-1\\ -1&0&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{!2}+\Delta_{11}\\ \Delta_{22}\\ \Delta_{21}\end{array}\right]$ (94)
Since $C1=F,C3=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{!2}+\Delta_{11}\neq 0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{21}\neq 0$ (95)
8.4.3 Sub-Case c $C1=F,C2=F,C3=T,C4=F$
Substituting $C3$ into $C1,$ $C2,$ $C4$ we have
$j$ $=$ $l+\Delta_{11}$
$l$ $=$ $k+\Delta_{12}$
$=$ $m+\Delta_{22}+\Delta_{12}$
$m$ $=$ $j+\Delta_{21}$ (96)
Thus we have the marix equation
$\left[\begin{array}[]{ccc}+1&-1&0\\ 0&+1&-1\\ -1&0&+1\end{array}\right]\;\left[\begin{array}[]{c}j\\ l\\ m\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{22}+\Delta_{12}\\ \Delta_{21}\end{array}\right]$ (97)
Since $C1=F,C2=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}\neq 0$ (98)
8.4.4 Sub-Case d $C1=F,C2=F,C3=F,C4=T$
Substituting $C4$ into $C1,$ $C2,$ $C3$ we have
$j$ $=$ $l+\Delta_{11}$
$l$ $=$ $k+\Delta_{12}$
$k$ $=$ $m+\Delta_{22}$ (99)
$=$ $j+\Delta_{21}+\Delta_{22}$
Thus we have the marix equation
$\left[\begin{array}[]{ccc}+1&0&-1\\ 0&-1&+1\\ -1&+1&0\end{array}\right]\;\left[\begin{array}[]{c}j\\ k\\ l\end{array}\right]=\left[\begin{array}[]{c}\Delta_{11}\\ \Delta_{12}\\ \Delta_{21}+\Delta_{22}\end{array}\right]$ (100)
Since $C1=F,C2=F,C4=F$ we require there not to be a solution.
As $|M|=0$ there is not a non-zero solution for $\mathbf{v}$ if the following requirements are satisfied
$\Delta_{11}\neq 0\quad And/Or\quad\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}+\Delta_{22}\neq 0$ (101)
8.4.5 All Sub-Cases of Case 4
Overall, we see that if the requirement is satisfied, we can say $\delta=1$.
Hence applying this result for $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(1\right)\,\right\}$ (102)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(1)$
$\leq$ 11 11
From above the following requirements must be satisfied for Case 4 to cover all sub-cases.
$\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{11}+\Delta_{21}\neq 0\qquad Sub-Case\text{ }a$
$\Delta_{11}+\Delta_{12}$$\neq$ $0\quad And/Or\quad\Delta_{22}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }b$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{22}+\Delta_{12}\neq 0\quad And/Or\quad\Delta_{21}\neq 0\qquad Sub-Case\text{ }c$
$\Delta_{11}$$\neq$ $0\quad And/Or\quad\Delta_{12}\neq 0\quad And/Or\quad\Delta_{22}+\Delta_{21}\neq 0\qquad\qquad Sub-Case\text{ }d$
The Bell inequality is then $S\leq 1$.